### v4

# Enhancing Efficiency in Multidevice Federated Learning through Data Selection

Abstract.
Federated learning (FL) in multidevice environments creates new opportunities to learn from a vast and diverse amount of private data. Although personal devices capture valuable data, their memory, computing, connectivity, and battery resources are often limited. Since deep neural networks (DNNs) are the typical machine learning models employed in FL, there are demands for integrating ubiquitous constrained devices into the training process of DNNs. In this paper, we develop an FL framework to incorporate on-device data selection on such constrained devices, which allows partition-based training of a DNN through collaboration between constrained devices and resourceful devices of the same client. Evaluations on five benchmark DNNs and six benchmark datasets across different modalities show that, on average, our framework achieves $\sim$19higher accuracy and $\sim$58lower latency; compared to the baseline FL without our implemented strategies. We demonstrate the effectiveness of our FL framework when dealing with imbalanced data, client participation heterogeneity, and various mobility patterns. As a benchmark for the community, our code is available at github.com/dr-bell/data-centric-federated-learning.
copyright: none
1. Introduction
Multidevice FL enables learning new insights from various clients by training DNNs on personal data in a more private manner (McMahan et al., 2017a; Kairouz et al., 2021). However, some wearable or embedded devices have limited computing resources, little on-device storage, and variable network connectivity. Most existing FL frameworks do not account for constraints of personal devices and instead assume either modern smartphones (Hard et al., 2018; Niu et al., 2020) or accelerators (Cho et al., 2022; Zhang et al., 2022) as target client devices. Although smartphones and accelerators are less powerful than cloud machines, they are nevertheless equipped with mobile GPUs, possess gigabytes of runtime memory, and have fairly stable connectivity to a central server, which simplifies the requirements of DNN training (Yang et al., 2021; Lin et al., 2020).
Challenge. We focus on multidevice FL for ubiquitous constrained devices (UCDs). Examples are wearable devices such as earbuds, glasses, or rings, and embedded devices such as environmental or industrial sensors. We are driven by the motivation that UCDs are becoming the primary data-producing devices for individuals and industries due to various on-device sensors, such as cameras and inertial measurement units. UCDs offer a wealth of spatio-temporal data that is frequently absent when focusing solely on smartphones or plugged-in devices. For instance, people often opt for wristbands, smart rings, or earbuds over carrying a smartphone in their hand or pocket while engaging in outdoor, industrial, or sports activities (Seneviratne et al., 2017; Mo and Zhou, 2021). Moreover, low-cost sensors distributed across farms or forests remotely monitor crops and livestock and detect potential threats (Papst et al., 2021). Such abundance and diversity in sensory data captured by UCDs can facilitate the learning of more effective models for emerging applications in personal and industrial environments. Nevertheless, there are two challenges for FL on UCDs:
The first challenge is the network connectivity to a central server. Usually, UCDs use an access point (AP) over wireless channels, e.g., a smartphone or a router, as a relay to connect to the central server. Such wireless connections, e.g., via Bluetooth Low Energy (BLE), are unreliable due to the devices’ mobility. Particularly for outdoor settings, UCDs may not always be connected to the companion AP, and hence therefore no immediate connection to the central server.
The second challenge emerges from constraints on the memory and compute resources available on UCDs, particularly the model size and functionality of DNNs that can be trained locally on UCDs. Although a naive solution is to transfer all data from UCDs to their respective APs and perform training of DNNs on APs, storing all data could exceed the storage capacity of UCDs, especially when UCDs disconnect from their AP often. To tackle these challenges effectively, we require FL frameworks that dynamically distribute data and computations between UCDs and APs.
Solution. We propose an end-to-end FL framework that orchestrates the local training of DNNs among UCDs and APs by integrating on-device data selection alongside partition-based model training. We refer to our proposed framework as Centaur. (A) Considering computation constraints, Centaur initializes the DNN in two partitions: an encoder (i.e., feature extractor) to be only trained on APs, followed by a lightweight classifier to be trained on both APs and UCDs (§3.2).
This is inspired by split (Gupta and Raskar, 2018; Thapa et al., 2022) and transfer (Weiss et al., 2016) learning, where UCDs keep the encoder frozen and only train the classifier, while the AP train the full DNN. (B) Considering memory and connectivity constraints, Centaur performs data selection (Jiang et al., 2019) by analyzing the training loss and the gradients norm of the classifier part, to decide which data points captured by UCDs contribute more to the training of which part of the DNN. Here, data points are categorized as either of (i) discarded if they have very low loss values, (ii) kept locally on the UCD to train the classifier part if their loss values or gradients norm are not very high, and (iii) transmitted to the AP to train both encoder and classifier part if they cause high values for both loss and gradients norm [^1] The assumption is each pair of UCD and AP belongs to the same client, thus transmitting data from a UCD to its AP does not violate the clients’ privacy. (§3.3).
Centaur differs from the prior works on hierarchical FL (Liu et al., 2020; Abad et al., 2020), since they neither perform partition-based training nor incorporate data selection. This makes our FL framework adaptive to the available resources in a multidevice setting.
Evaluation. We perform evaluations on five benchmark DNNs and six benchmark datasets. We show that compared to existing FL alternatives, Centaur not only saves more bandwidth by reducing the communication cost associated with offloading samples from the UCD to the AP, but also reduces the latency of training, and as a result, the energy consumption, on both the AP and the UCD. As a prime example of our experimental results, Figure 1 shows the latency vs. accuracy of performing FL for two different tasks: (1) four benchmark DNNs, with their encoder pre-trained on ImageNet (Deng et al., 2009) , and then finally trained for CIFAR10 (Krizhevsky et al., 2009) image classification with 100 clients, and (2) a benchmark ConvNet (Chang et al., 2020), trained from scratch, on three human-activity recognition (HAR) datasets with 20 clients. In both figures, we compare Centaur with the baseline of standard FL on UCDs without our implemented strategies. Centaur, through data selection and partition-based model training, achieves up to $19\%$ higher accuracy and $58\%$ lower latency, on average. We provide thorough experiments and comparisons with other alternative baselines in §6.
Figure 1. Training latency vs. test accuracy of Centaur compared with standard UCD training (i.e., only training the classifier on UCDs without data selection and partition-based training). (Left) Four benchmark pre-trained DNNs fine-tuned on CIFAR10 dataset with FL. (Right) A ConvNet trained with FL on three benchmark wearable sensor datasets: MotionSense, PAMAP2, and UCIHAR.
Contribution. Our solution enables UCDs to efficiently participate in multidevice FL alongside their upstream APs, by incorporating a customizable data selection scheme and a partition-based training regime. The impact is a decrease in computational and communication costs, as well as an improvement in the model’s test accuracy. Our experiments on a small testbed of RaspberryPis validate that Centaur is promising in decreasing overall storage requirements and training time due to employing effective data selection. Moreover, our analysis on the mobility of UCDs accounts for the real-world spatiotemporal mobility of edge devices participating in FL. Our extensive empirical analyses, including measurement of cost, imbalance data, participation heterogeneity, and connection probability, indicate that Centaur achieves higher efficiency (i.e., high accuracy and low cost) across most scenarios. Our code is open-source to encourage further advancements in this domain.
2. Identifying Design Challenges: A Motivational Study
Terminology. A ubiquitous constrained device (UCD) is a mobile device with restricted hardware and connectivity resources. An access point (AP) is a resourceful device that has much better hardware and connectivity capabilities than its corresponding UCD. A server is a central entity owned by a service provider that organizes FL over several clients. An encoder ($E$) is the first part of the DNN and extracts the features of input data. A classifier ($C$) is the second part of the DNN and classifies the encoded features into a final prediction. Let $\mathbf{X}$ denote the input data and $\mathbf{y}$ denote its label.
2.1. The Study
We perform measurement experiments to identify the key design requirements that can facilitate the deployment of FL on constrained devices.
Setup. We create a testbed with four RaspberryPi 4 Model B as the clients (Velasco-Montero et al., 2018; Drakopoulos et al., 2019); all running the 64-bit Raspbian OS with total secondary storage of $32$GB. We set up a cohort of clients having varying primary memories, with clients C1, C2, C3, and C4 having $1$GB, $2$GB, $4$GB, and $2$GB of primary memory, respectively. For analyzing the CPU and memory consumption of the clients, we use the logs from the /proc/stat and free system calls, respectively. For the model architecture, we use a benchmark architecture MobileNetV3 (Howard et al., 2019) (pre-trained on ImageNet) as the encoder followed by a medium-sized classifier with two fully connected layers with an intermediate drop-out layer. We use CIFAR10 dataset and measure the resource and time consumed on-device for both training, per each epoch, and inference. Finally, for implementing FL setup, we use Flower (Beutel et al., 2022) over four RaspberryPi devices and a Mac Mini as the server, as shown in Figure 2a.
Running Standard FL. We implement standard FL (for $100$ rounds) over four RaspberryPi clients with the standard FedAvg (McMahan et al., 2017b). We consider two scenarios: (1) when only the classifier part is trained, and (2) when the entire model (encoder and classifier) is trained. For the first scenario, we observed in our experiments that the standard FL runs successfully over all four clients. However, for the second scenario, client C1 with 1GB of primary memory fails to participate in the FL process. Notably, 1GB of primary memory is still significant in terms of memory capability and typical UCDs usually have a much lesser amount of memory than this. Therefore, it is clear that in such cases, it is much more challenging to deploy standard FL over UCDs. To further assess this, we conduct a series of experiments in the following section to investigate the resource and time consumption for training DNNs on our testbed.
On-device Training. Considering the challenges discussed previously, we analyze the latency and resource consumption for local training of MobileNetV3. Figure 2 shows that there is a significant rise in the overall CPU and memory consumption when the entire model (encoder and classifier) is trained on-device in comparison to when only the classifier is trained and the encoder kept frozen. A key observation is that by training only the classifier, we require less processing power and approximately the same amount of memory resources as when the entire model is used solely for inference. Moreover, Figure 3 shows that the time taken to train the classifier is less than the time to perform a single forward pass on the entire model. These measurements confirm that training the classifier only is 50×50$\times$50 × time efficient compared to training the entire model, which helps in reducing the battery consumption significantly. When we follow a naive approach of randomly dropping samples from the training dataset, we observe a decline in the overall training time of the classifier (see Figure LABEL:fig:select_train). However, this comes at the cost of reduced accuracy, which might be detrimental to the overall system.
Figure 2. Resource consumption of training DNNs on (a) the designed setup over four RaspberryPi 4 Model B. We compare (b) CPU Consumption and (c) Memory Consumption for three scenarios when: (i) only the classifier is updated, (ii) the whole model is updated, and (ii) the whole model is used only at inference mode.
2.2. Insights
Here, we summarize the understandings derived from our motivational study:
(1) Resource and Training Efficiency. As UCDs typically have limited memory, training on the device should be memory efficient. This motivates using a frozen pre-trained encoder, and only training the classifier part on the device (Mathur et al., 2021). However, freezing the weights of the encoder limits the expressiveness of the model and prevents the model from learning important features of the newly collected data over the UCDs. Thus, while minimizing the on-device model training on the UCDs, our system should ensure that the encoder is periodically trained on a portion of new data collected from the UCDs, potentially on a more resourceful device where the entire model can be trained.
(2) Data Efficiency and Spatiotemporal Coverage. Typically, UCDs are deployed to maximize the spatiotemporal coverage of activities and events. Hence, our design should leverage the spatiotemporal richness of such data to train a better model. Such spatiotemporal diversity also makes UCDs more prone to intermittent connectivity, which may require the framework to store training data on-device. UCDs possess limited secondary storage, and even when they have persistent connectivity to the AP, streaming high volumes of data could drain their battery. Thus, our design should dynamically select a small subset of data points for storing locally or transmitting to the APs. This needs to be designed in a sophisticated manner such that it does not compromise the overall performance of the system.
Figure 3. (a) Time for backpropagation across (i) the entire model (encoder and classifier), (ii) across the classifier only, and (iii) for forward pass on the entire model. (b) The trend of training time and accuracy (for one epoch) for different sizes of training sets.
3. Our Proposed Framework: Centaur
With our understanding developed from the motivational study, we introduce our proposed framework. Figure 4 gives a schematic overview, and below we elaborate on the algorithms used in Centaur.
3.1. The Overview
The first level includes UCDs, which are connected to their APs via BLE, Wi-Fi, Backscatter, or other wireless technologies. We assume the connection between UCD and AP is not always available or reliable due to the natural mobility of UCDs. We consider that the data selection procedure and the training of the classifier only happen on UCDs. The second level, which includes APs, is responsible for training the encoder in conjunction with the classifier atop it. The third level includes a central server, which performs model aggregation for both classifiers and encoders that are trained by UCDs and APs. Here, we present a step-by-step description of our proposed framework.
Step \⃝raisebox{-0.9pt}{1}: We assume that the server receives the information about UCDs’ resource. Accordingly, the server initializes a DNN model $M$ that is formed of an encoder $E$ and a classifier $C$, and then distributes the $E$ and $C$ to the APs and subsequently to UCDs. We explain model initialization details in §3.2.
Step \⃝raisebox{-0.9pt}{2}: UCDs start FL by performing data selection on the collected data $D$, based on the computed loss value via the forward pass and the magnitude of the gradient of the classifier’s last layer. Each sample point will be categorized in one of (i) $D_{N}$ to be discarded, (ii) $D_{C}$ for locally training $C$ on the UCD, or (iii) $D_{M}$ to be transmitted to AP for training $M$. We explain data selection details in §3.3.
Step \⃝raisebox{-0.9pt}{3}: The UCD discards data $D_{N}$ in the current epoch without any further training steps or backward pass on them.
Step \⃝raisebox{-0.9pt}{4}: The UCD performs backward pass through the classifier $C$ on data $D_{C}$, while the encoder $E$ is frozen. Then, the UCD updates the parameters of $C$ via computed gradients. Notice that, in practice, the gradient computation done in Step \⃝raisebox{-0.9pt}{2} can be reused for this update.
Step \⃝raisebox{-0.9pt}{5}: UCDs share their updated classifier $C$ to the server, and the server performs aggregation on all $C$ from participating UCDs to obtain one aggregated classifier.
Step \⃝raisebox{-0.9pt}{6}: The server sends aggregated classifier back to the UCDs and APs.
Step \⃝raisebox{-0.9pt}{7}: With the updated classifier, the AP trains the encoder $E$ together with the classifier $C$ atop $E$ on data $D_{M}$. That is, $D_{M}$ is fed for a forward pass and then a backward pass throughout to update $M$.
Step \⃝raisebox{-0.9pt}{8}: The updated $M$ (including $E$ and $C$) is shared from the AP to the server. After collecting updated $M$ from all participating APs, the server performs model aggregation to obtain the aggregated encoder $E$ and classifier $C$ and distribute them back to UCDs and APs as in Step \⃝raisebox{-0.9pt}{6}. See § 3.4 for more details of FL training procedures. Steps from \⃝raisebox{-0.9pt}{2} to \⃝raisebox{-0.9pt}{8} will be repeated for several FL rounds to update $C$ on UCDs and $E$ on APs.
Figure 4. The overview of our framework, including model initialization (§ 3.2), data selection (§ 3.3), and partition-based training and aggregation (§ 3.4). We explain the set up in § 3.1 and elaborate the details of Steps \⃝raisebox{-0.9pt}{1} to \⃝raisebox{-0.9pt}{8}.
3.2. Model Initialization
To partition a DNN into an encoder $E$ and a classifier $C$, one option is to use pre-trained benchmark DNNs; considering numerous well-trained models, e.g., in Model Zoo (Koh, 2019). Usually, DNN architectures from a candidate space of pre-trained models have already been well developed, with efficient quantization and compression capabilities, to be deployed on UCDs for inference (Fedorov et al., 2019; Banbury et al., 2021; Xu et al., 2022; Liberis et al., 2021; Malekzadeh et al., 2021; Li et al., 2021a). We aim to utilize the full capability of UCDs for on-device training of the classifier $C$. Thus, we need to design the classifier such that it can run on the limited memory and computation power available on these devices.
To achieve these, we can iteratively look into candidate architectures that fit in the limited memory available on these UCDs. A robust mechanism in this direction is to obtain a history of memory usage of a particular device which reflects the typical memory availability on the UCD. Formally speaking, the iterative process can be viewed as an optimization problem described as
(1) $C=\min\{m_{1},m_{2},\ldots,m_{k}|\mathbb{P}(m_{k})\times\mathbb{B}<\mathbb{M}\},$
where $\mathbb{P}(m_{k})$ provides the number of parameters of the candidate architecture $m_{k}$, $\mathbb{B}$ is the number of bytes used to store the parameters and intermediate results in computations, and $\mathbb{M}$ is the memory available in the device for FL participation. The parameter $\mathbb{M}$ can be measured and logged as empirical observations of the UCD memory usage, to be sent to the server as we discussed in Step \⃝raisebox{-0.9pt}{1} in Figure 4.
In this paper, we limit our classifier architectures with fully connected layers; thus, we find a solution by either reducing the number of neurons or the layers to find a potential architecture for $C$ that can fit into the requirements for $\mathbb{M}$ in Equation (1). In the same vein, we tackle the problem of low computing resources by choosing the smallest architecture for $C$ and then deploying it on the UCDs.
3.3. Data Selection
We determine the importance of each data sample, along the forward pass, to decide whether to further perform (the more costly) backward pass on this sample or not. Such data selection satisfies design requirements by improving both storage efficiency and training efficiency. We use a combination of (i) the loss value and (ii) the norm of last-layer gradients to measure the sample’s importance.
3.3.1. Loss-based Selection.
All available data $D$ at round $r$ are fed into the model $M$ to compute their loss values $\ell_{r}(f_{M}(\mathbf{X}),\mathbf{y})$. We consider lower loss values as an indication of being less important, and higher loss values show the importance of the data (Jiang et al., 2019).
[…]
(2) $\begin{cases}\mathcal{P}_{N}(\ell_{r+1})=1-[\text{CDF}^{\ell}_{r}(\ell_{r+1})]^{\alpha}\\ \mathcal{P}_{M}(\ell_{r+1})=[\text{CDF}^{\ell}_{r}(\ell_{r+1})]^{\beta}\end{cases}$
where $\alpha$ and $\beta$ parameters determine the level of selectivity. At round $r+1$, for a sample with loss value $\ell_{r+1}$, the sample is selected into $D_{N}$ with probability of $\mathcal{P}_{N}(\ell_{r+1})$. This implies that the lower the loss value for a sample, the greater the probability it will be discarded. Similarly, $\mathcal{P}_{M}$ is used to select samples with the highest losses as $D_{M}$; which means that the higher the loss value for a sample, the greater the probability it will be chosen for training the entire model.
[…]
3.3.2. Gradient-based Selection.
[…]
(3) $\mathcal{P}_{M}^{+}(\|g_{r+1}\|)=[\text{CDF}^{g}_{r}(\|g_{r+1}\|)]^{\gamma}$
where $\gamma$ is used for customizing the selection rate. Samples with high norm are selected with $\mathcal{P}_{M}^{+}(\|g_{r+1}\|)$ and then added to data $D_{M}$. Our dynamic strategy, in combining loss and last-layer gradient norm, enables Centaur to achieve a better trade-off between computation cost and selection performance on UCDs. Moreover, one may substitute this module with a different data selection technique that could also be dependent upon the specific use case.
321 Inputs: (1) $\mathcal{A}$: the set of $A$ clients each having two devices, one UCD and one AP, (2) $D^{k}$: local dataset for each client $a\in\{1,\dots,A\}$, (2) $f_{E}$: the encoder part with parameters $W_{E}$, (3) $f_{C}$: the classifier part with parameters $W_{C}$, (4) $R$: the total number of FL rounds. for
$r\in\{1,\dots,\nicefrac{{R}}{{2}}\}$ rounds
 do
6      54 $K$ clients $\Leftarrow$ randomly select $K$ clients from a total of $A$, using uniform sampling — Training on UCDs — for $k\in\{1,...,K\}$ do
11            10987 $\mathbf{X}^{k},\mathbf{y}^{k}\leftarrow D^{k}$ (all the samples in the local dataset of client $k$ UCD) $\ell^{k}=\mathcal{L}(f_{E}\circ f_{C}(\mathbf{X}^{k}),\mathbf{y}^{k})$ (forward pass to compute per-sample loss values) $D_{N}^{k},D_{C}^{k},D_{M}^{k}\Leftarrow$ based on $\ell^{k}$, perform data selection as detailed in Section 3.3 for all $(\mathbf{X}^{k}_{j},\mathbf{y}^{k}_{j})\in D_{C}^{k}$ do
16                  15141312 $g^{k}_{C}={\partial\ell^{k}}/{\partial W_{C}}$ (compute $f_{C}$’s gradients on $D_{C}^{k}$ to update $f_{C}$)$W_{C}^{k}=Optimizer(W_{C}^{k},g^{k}_{C})$ (update $f_{C}$’s parameters using gradients) $D^{+}\Leftarrow$ based on $g^{k}_{C}$, perform data selection as detailed in Section 3.3 $D_{M}^{k}=D_{M}^{k}\cup D^{+}$ (enhance $D_{M}^{k}$ by high-value gradients data points)
20      191817 — Server Side Aggregation on the Classifier Part — $W_{C}\leftarrow\frac{1}{K}\sum_{k\in K}W_{C}^{k}$ $W_{C}^{k}\leftarrow W_{C}$ (update all clients with new $W_{C}$) — Training on APs — for $k\in\{1,\dots,K\}$ do
25            24232221 $\mathbf{X}^{k},\mathbf{y}^{k}\leftarrow D_{M}^{k}$ (all selected samples for training $f_{E}$ on client $k$ AP) $\ell^{k}=\mathcal{L}(f_{E}\circ f_{C}(\mathbf{X}^{k}),\mathbf{y}^{k})$ (forward pass to compute per-sample loss values) $g^{k}={\partial\ell^{k}}/{\partial(W_{E},W_{C})}$ (compute gradients on $D_{M}^{k}$ to update $f_{E}$ and $f_{C}$) $W_{E}^{k},W_{C}^{k}=Optimizer(W_{E}^{k},W_{C}^{k},g^{k})$ (update all parameters using gradients)
28      2726 — Server Side Aggregation on both Encoder and Classifier Parts — $\{W_{E},W_{C}\}\leftarrow\frac{1}{K}\sum_{k\in K}\{W_{E}^{k},W_{C}^{k}\}$$\{W_{E}^{k},W_{C}^{k}\}\leftarrow\{W_{E},W_{C}\}$ (update all clients with the new aggregated model)
29 return $\{W_{E},W_{C}\}$
Algorithm 1 Multidevice FL with Data Selection and Partition-based Training
3.4. Partition-based Training and Aggregation
The training process is conducted on both UCDs and APs. While UCDs only train the classifier $C$ on data $D_{C}$, APs train the complete model $M$ on data $D_{M}$. Such procedure allows the model to be (partially) updated when UCDs are offline, in addition to the full updates when UCDs have a connection to the internet. This better utilizes UCDs’ spatiotemporal richness and consequently improves the training efficiency in our design requirements.
Algorithm 1 shows the process of partition-based training and aggregation in Centaur. We set the algorithm to run $\nicefrac{{R}}{{2}}$ rounds, considering that UCD training and AP training proceed iteratively in one round. All training is conducted on the client’s devices either on UCDs or APs. Then with all updated weights, the server performs Federated Averaging (FedAvg) (McMahan et al., 2017b) to obtain one aggregated global model. The new global model needs to be distributed to all clients UCDs and APs such that (i) the training on AP has an updated classifier, and (ii) the training on UCD in the next round has both an updated encoder and classifier.
4. Metrics
To evaluate the performance of our FL framework, compared to other FL alternatives, we use the following metrics:
1) Accuracy. This is the classification accuracy of the trained model on a test set hosted by the central server. For a model obtained at the end of each training round, the test accuracy is computed as
$acc\equiv\frac{1}{T}\sum_{i=1}^{T}\mathbb{1}\big{(}\operatorname*{argmax}\big{(}f_{E}\circ f_{C}(\mathbf{X}_{i})\big{)}=\mathbf{y}_{i}\big{)},$
where test set has $T$ pairs of $(\mathbf{X}_{i},\mathbf{y}_{i})$ and $\mathbb{1}(\mathrm{C})$ denote the indicator function that outputs 1 if condition $\mathrm{C}$ holds.
2) Multiply–Accumulate (MAC). This operation is a common step that computes the product of two numbers and adds that product to an accumulator ($a\leftarrow a+(b\times c)$); a fundamental operation for both any DNN layers during training and inference. We use fvcore (FAIR, 2019) library to compute the number of MAC operations. Since fvcore only supports counting MACs in a forward pass, and to count MACs in a backward pass, we use the heuristic that FLOPs (i.e., double MACs) ratio of the backward-forward pass is typically between 1$\times$ and 3$\times$ and most often is 2$\times$ based on models’ specific layer types, according to previous observations (Marius Hobbhahn, 2021).
3) Bandwidth. As the model size (i.e., the encoder or the classifier) can be different in each FL round, we use fvcore to count the number of parameters that are communicated in each round. Based on the model size and number of communications among UCDs, APs, and the server, we compute the amount of bandwidth that is consumed. When it is a case, we also count the number of sample points that are uploaded from UCDs to APs.
4) Latency. The latency usually has a linear relation with MAC operations; due to the lack of specialized accelerators (Liberis et al., 2021; Lin et al., 2022). Thus, we estimate the latency of model training based on the processor’s frequency:
(4) $\text{Latency}=c*\frac{\text{Total MACs}}{\text{Processor Frequency}}.$
The ratio between MAC operations and the processor’s instructions, $c$, is typically between $1$ and $2$ based on specific instruction sets/compilers. For simplicity, we assume that each MAC operation translates to two instructions in an MCU (i.e., $c=2$). Notice that existing processors (e.g., Intel’s Load Effective Address) completes one MAC in one instruction (WiKiBooks, 2021)), and this only scales the experimental results and does not change the final conclusions. Finally, the latency of communication is estimated based on the total amount of data needed to be transmitted, divided by the up-link speed or the down-link speed of devices.
5) Energy. This is the total execution time multiplied by the processor’s consumed power per unit of time. The energy consumption of communication can also be calculated based on the total time of transmitting data multiplied by the transmitters’ power per unit of time. We set the values for UCDs and APs in our simulations as follows:
(A) As computation resources, we set up:
Device
CPU Frequency (MHz)
Storage Capacity (MB)
Power Capacity (mW/MHz)
Timeout (s)
UCD 100 5 0.05 Inf
AP 2000 4096 1.5 Inf
(B) For communication resources, we set up:
Device
Uplink(Mbit/s)
Downlink (Mbit/s)
Communication Energy (W)
Disconnect Probability
UCD 2 2 0.0001 0.5
AP 10 100 10 0
5. Experimental Setup
Datasets. We use six commonly used datasets. CIFAR10 and CIFAR100 (Krizhevsky et al., 2009) that both include 50K training samples and 10K test samples, with 10 and 100 classes respectively. EMNIST (Cohen et al., 2017) that includes 112K training samples with 47 classes. UCIHAR (Anguita et al., 2013) is a widely used dataset of 30 users performing 6 daily activities; data from the accelerometer and gyroscope sensors were collected by a smartphone worn on the waist. Data from 21 users are used for training and that of the other 9 users for testing purposes. MotionSense (Malekzadeh et al., 2019) also includes accelerometer and gyroscope data from 24 users with a smartphone in the pocket of the trousers who performed 6 activities in 15 trials. We use as test data one trial session for each user and as training data the remaining trial sessions (e.g., one trial of “walking” of each user is used as test data and the other two trials are used as training). PAMAP2 (Reiss and Stricker, 2012) dataset contains data of 13 different physical activities, performed by 9 subjects wearing 3 devices and a heart rate monitor. We use the training-test split provided by PersonalizedFL library (Lu and Wang, 2023).
Models. For image classification datasets, we select four DNNs commonly used in mobile/edge-oriented literature: (1) EfficientNet-v2 (Tan and Le, 2021), (2) MobileNet-v3 (Howard et al., 2019), (3) MNASNet (Tan et al., 2019), and (4) ShuffleNet (Ma et al., 2018), with 5.3M, 2.2M, 2.5M, and 1.4M number of parameters respectively. For human-activity recognition datasets, we borrow the ConvNet architecture proposed in (Chang et al., 2020). All these models consist of an encoder made of multi-layer CNNs followed by one classifier.
FL Settings. We use Flower (Beutel et al., 2020), a customizable open-source FL framework (Python v3.7, Ray v1.11, and Torch v1.12). We run all simulations on a server with 80 Intel Xeon(R) E5-2698 CPUs, 8 Tesla V100 GPUs (16GB), and 504GB system RAM. For all simulations, we use 100 communication rounds, where at each round, UCDs and APs successively train the model. Unless specified otherwise, we consider 100 clients (each having one AP and one UCD). We randomly sample 10% of the clients for training in each round in our initial experiments and later show that our results hold when the client participation ratio is gradually increased to 100%. Data ($D$) are partitioned using Latent Dirichlet Allocation (LDA) (Blei et al., 2003; Beutel et al., 2022) without resampling (LDA-Alpha=1000). We refer to prior research (Blei et al., 2003) about details of LDA’s generative process. After such partition, each client owns $\nicefrac{{|D|}}{{100}}$ local samples.
Figure 5. Test accuracy when using different values for parameters $\alpha$, $\beta$, $\gamma$ in data selection.
[…]
Baselines. (1) AP Training. In this baseline approach we consider UCDs as data collection apparatuses only and therefore, UCDs upload collected and stored data samples to their connected APs, so that the training only happens on APs. In this case, the complete model (both the encoder and classifier) is updated because APs in general are considered to have sufficient resources. (2) UCD Training. UCDs do not upload the collected data samples. Instead, they conduct training on the data locally. However, as UCDs are typically resource-constrained, only the classifier is considered to be trained on-device. This means that the encoder part is considered to be frozen in this setting.
Mobility Model. In our case, the idea of the mobility model signifies how the mobility patterns of the UCDs (or users) impact the connectivity of the UCD to its AP and eventually to the internet. We define an exclusive-Online Association Matrix $\Omega$, which is a binary matrix representing the user’s exclusive location at a given time instance. Mathematically, $\Omega$ is a binary matrix $\mathcal{T}\times\mathcal{S}$ where $\mathcal{T}$ and $\mathcal{S}$ represent the temporal and spatial granularity of $\Omega$ marix, respectively. For example, the rows may represent the time zones of the day (early morning, morning, afternoon, evening, and late night), during which the user moves between three locations like home, office, and a public park, represented by the columns of $\Omega$. Furthermore, we ensure that each row’s sum equals unity, which means the user is present exclusively at a unique location in a given temporal instance. Finally, to simulate the connectivity patterns, we define a connectivity matrix $\lambda$, a single row-vector of dimension $\mathcal{S}$, denoting the connectivity probability across different locations. More specifically, for evaluating Centaur during mobility, we generate the global connectivity matrix $\lambda$ and the generate matrix $\Omega$, both chosen uniformly at random, for each user to simulate the mobility scenario.
6. Evaluation Results
We present the results when running Centaur compared to conventional FL training methods. We also analyze the impact of data/participation heterogeneity and spatiotemporal coverage on training methods.
Figure 6. Accuracy of the best classifier for four different encoders, trained on CIFAR10, CIFAR100, and EMNIST. Dash lines ( ) above depict the upper-bound accuracy when UCD devices have no resource and connectivity constraints.
6.1. Model Accuracy
We use the CNN backbone of the four benchmark models described in §5 as the encoders. We examine three classifiers: small that is only one fully connected (FC) layer of size $z$ (number of classes), medium that has two FC layers of size 64 and $z$, and large that has two FC layers of size 128 and $z$.
Centaur outperforms AP training and UCD training for different encoders. We permute the four encoders and three classifiers and report the test accuracy of Centaur compared to the two other baselines: AP training and UCD training. In Figure 6 we report the highest test accuracy for each encoder across all three classifiers for three datasets. Results show that Centaur outperforms AP training by $0.53\%\sim 40.15\%$ and UCD training by $0.45\%\sim 17.78\%$, depending on the settings. There are several settings in which all FL training methods cannot reach a good accuracy, e.g., NASNet and ShuffleNet on CIFAR100, probably because of their relatively small model sizes compared to data complexity. Among them, Centaur still achieves better performance.
Centaur outperforms UCD training for different classifiers. In Figure 7 we present the accuracy of different classifiers for each encoder. In the top plot, the results of training on CIFAR10 and CIFAR100 with an accuracy higher than $30\%$ are presented for better visibility. In the bottom plot, we report the results of UCIHAR, MotionSense, and PAMAP2. The results show that Centaur outperforms both UCD training in all classifiers’ sizes. The test accuracy also tends to be similar across small, medium, and large classifiers. In addition, the classifier’s sizes may have less impact on more sophisticated encoders (e.g., EfficientNet and MobileNet), which is also observed in a previous work (Mo et al., 2021). This may be because an appropriate encoder already produces high-quality features for unseen data that are easy to learn (e.g., CIFAR10), and in such a case, classifier sizes do not make any difference in test accuracy. Similar patterns are observed in HAR datasets. Notice that for HAR datasets, there is currently a lack of publicly accessible pre-trained encoders to start FL with. As a result, we can observe a more significant performance gap compared to image classification, particularly for more complex datasets, e.g., PAMAP2. This lack of pre-trained encoders can also be the reason that, in HAR datasets, the performance of Centaur and AP are almost the same.
Figure 7. Centaur performance vs. other baselines for different sizes of classifiers: small, medium, and large. (Top) image classification, and (bottom) HAR datasets.
Figure 8. Computation workload of model training on usage-constrained devices (Top Left) and access points (Top Right), and Communication workload of model/data sharing of usage-constrained devices (Bottom Left) and access points (Bottom Right) when achieving specific test accuracy, reported on MobileNetv3 and CIFAR10.
6.2. Efficiency
Cost-Accuracy Trade-off. We use CIFAR10, and fix MobileNet-v3 as the encoder and medium size for the classifier. We compute the test accuracy for a range of MAC, Bandwidth, Latency, and Energy budgets. The average size of each sample of CIFAR10 is $30KB$. As shown in Figure 8 (top), on both UCD and AP and for all ranges of accuracy, Centaur achieves lower MAC and Latency than both UCD training and AP training. In addition, Figure 8 (bottom) shows that communication Bandwidth and Latency of Centaur is almost the same as AP training; however Centaur can achieve a higher test accuracy along with more training steps while consuming more resources. Centaur causes much more communication cost than UCD training, because the training only involves the transmission of the classifier, which is much lighter than the encoder and data samples transmitted by Centaur. However, the accuracy of UCD training cannot go further than $83.10\%$, while Centaur can reach up to $89.90\%$ accuracy. We remark that model training consumes much more energy and causes much more latency than communication.
Correlation among Cost Metrics. To show the connection between cost metrics, we plot MAC, Latency, and Energy in Figure 9. We observe that both Latency and Energy have linear relationships with MAC. Besides, Latency is also linearly correlated with Energy, in such a way that these two entirely overlap when we appropriately scale the y-axes. We remark that Bandwidth also has a similar correlation. Such linear relationships are because of the assumption we made when computing Latency and Energy; however, the actual cost and the relationship among these metrics might not deviate much in practice.
Based on the above-mentioned correlation, we can add Latency (or Energy) of model training as well as the communication Bandwidth to get the overall workload. The results shown in Figure 1 indicate that the workload still follows a similar pattern as the computation in model training, because the training workload significantly overweight the communication workload.
6.3. Data and Participation Heterogeneity
Centaur is robust to data heterogeneity. To create imbalanced non-IID data partitions among clients, we use LDA as defined in § 4. Next, we set different values for LDA-Alpha to manipulate the levels of non-IID data partitions. Figure 11 and Figure 11 show the results when LDA-Alpha are changing from $0.001$, to $0.1$, to $1$, and to $1000$. The smaller is LDA-Alpha, the less balanced is the dataset. We also annotate some examples of class distribution in Figure 11 demonstrating that LDA-Alpha$=0.001$ generates almost $1$ class per client, while LDA-Alpha$=1000$ generates almost uniform classes per client. Results show that in all cases Centaur can reach higher test accuracy than both AP and UCD training. We remark that due to the retraining of the encoder, both Centaur and AP training can cause more fluctuations than UCD training; this is especially obvious when data distribution tends to be IID (e.g., when LDA-Alpha$=1000$ in Figure 11).
Centaur scales well with higher total and participating clients. We further scale the FL training problem with the number of total clients from 10 to 1000. $10\%$, $20\%$, and $50\%$ of them are selected respectively as participating clients in each round. Figure 12 shows Centaur achieves higher test accuracy compared to both UCD training and AP training. Specifically, with a higher participation rate, all training methods tend to have better performance. However, with a much larger number of total clients (e.g., 1000), the test accuracy is reduced, because the same size of data is partitioned into more portions. Also, with such more fragmented data partitioning, Centaur has much accuracy gain when compared to UCD training.
Figure 9. Cost linkage analysis among MAC operations, latency, and energy consumption.
Figure 10. The impacts of imbalanced data partitioning on the test accuracy of AP training, UCD training, and Centaur, reported on MobileNetv3 on CIFAR10 dataset (annotated texts give examples of the classes present on the UCDs)
Figure 10. The impacts of imbalanced data partitioning on the test accuracy of AP training, UCD training, and Centaur, reported on MobileNetv3 on CIFAR10 dataset (annotated texts give examples of the classes present on the UCDs)
Figure 11. Impacts of imbalanced data partitioning on the test accuracy of AP training, UCD training, and Centaur, reported on ConvNet and UCIHAR dataset dataset.
6.4. Performance of Centaur under Mobility
We evaluate situations when the connection probability $\lambda$ changes based on the mobility model defined in §4. For a more readable visualization, we bucket $\lambda$ in three different ranges of $\lambda\in$ [0.1,0.4], $\lambda\in$ [0.4,0.7], and $\lambda\in$ [0.7,1.0] and report the corresponding Accuracy of Centaur when the UCDs are mobile. Results reported in Figure 13 show that Centaur always has higher efficiency than standard UCD training while incurring a higher latency than direct AP training. Specifically, regarding the much important UCD side, Centaur gains $6.45\%$, $5.64\%$, and $4.11\%$ higher accuracy with $51.73\%$, $60.48\%$, and $68.35\%$ lower cost both in terms of energy and latency across the different ranges of $\lambda$.
Interestingly, we also observe a rise in accuracy with lower connectivity probability. The primary reason is attributed to the overall design of our experimental setup whereby we make a realistic assumption that when mobile with limited connectivity, these UCDs can gather more data and perform the data selection. Additionally, with this judiciously selected data, the device trains the classifier for an extended number of epochs thus boosting its accuracy.
7. Discussion and Related Work
Multitier FL. Compared to the default client-server FL, client-edge-server FL (Liu et al., 2020; Abad et al., 2020) considers a middle layer, e.g., a cellular base station, to orchestrate FL training across clients and a central server. In such a setup, clients in the last tier participate with several devices (Cho et al., 2022; Diao et al., 2020; Zhang et al., 2021). This is shown to achieve better trade-offs between accuracy and bandwidth consumption. However, these works assume local training of the whole model on the client devices and do not utilize the capabilities of a resourceful edge device for aiding a constrained device during training, as they are only used for aggregation/communication. This makes it impractical in scenarios involving UCDs.
On-device Training. Most prior works only consider the challenges of on-device inference (i.e., forward pass) (Fedorov et al., 2019; Banbury et al., 2021; Xu et al., 2022; Liberis et al., 2021), or work on sparse training (Li et al., 2021a; Bibikar et al., 2022). Sparse training or pruning still requires computational resources (especially memory) that UCDs do not have. Few works propose solutions for full on-device training (i.e., forward and backward pass) (Lin et al., 2022), but not for FL scenarios as they often make impractical assumptions like the availability of large datasets for network architecture search (Liberis et al., 2021) or unconstrained memory for hosting the full model (Diao et al., 2020; Zhang et al., 2021). Transfer learning can be used on UCDs to fix the encoder architecture, which in turn restricts the extraction of newer features, and only trains the last layer for the personalized training (Zhang et al., 2021). Regarding complex tasks, training the fully functional model on UCDs is challenging, because the on-device model still needs the capability of learning all features by itself (Li et al., 2021a; Bibikar et al., 2022), and the computational resources at the upstream devices are unutilized.
Figure 12. The impact of participating clients and total clients on test accuracy reported on MobileNetv3 and CIFAR10.
Communication/Power Efficiency. Using a pre-trained encoder without further training might fail to capture discerning features from the newly generated data, causing lower performance for the ultimate classifiers. On the other hand, passing all collected sensor data from UCDs to APs (even with data compression) is not always feasible and it might consume a lot of energy and bandwidth. Importantly, network availability for UCDs can be fragile due to the low-energy wireless communication protocols they use and the mobility of these UCDs. As UCDs are becoming stand-alone, it is expected that they can operate as independent devices without internet or permanent connection to an upstream device, Centaur takes a major step in unlocking more functionalities of such development. Compression techniques (Lim et al., 2019) or topology-aware hierarchical aggregation (Luo et al., 2020) can reduce Latency by taking advantage of the locality in networks and continuously balancing workloads across clients. However, compression and hierarchical aggregation alone do not allow for seamlessly incorporating UCDs in FL. Our proposed method comprises different methods to offer customizable trade-offs between test accuracy and communication cost in data- and topology-agnostic manners, while allowing additional use of the recent compression and hierarchical aggregation techniques for further efficiency.
Training on UCDs. By default, FL algorithms are not specifically designed for a multi-device environment (Lee et al., 2019; Islam and Nirjon, 2019). Training DNNs on tiny MCU devices is extremely challenging due to limited memory, computation, and power resources. There are recent efforts to enable running ML models on a micro-controller with highly limited memories, such as TensorFlow Lite Micro (David et al., 2021) or Tiny Training Engine (Lin et al., 2022). Such developments, usually based on a combination of hardware and software novelties, can be easily integrated into Centaur to provide further efficiency. As we observe, a better training framework can not only enable deploying larger classifiers on UCDs, but can also allow us to calculate more gradients for better data selection. While it is difficult to train large DNNs on UCDs, these devices are typically parts of a personal network, and therefore, they can privately communicate/coordinate with other surrounding devices. Our work considers such a hierarchy for unlocking more utilities while providing better efficiency.
Figure 13. The performance of access point (AP) training, usage-constrained device (UCD) training, and partition-based federated learning (Centaur) under different connection probabilities $\lambda$, reported on MobileNetv3 and CIFAR10.
8. Limitations and Future Work
First, we accommodate FedAvg (McMahan et al., 2017b) in our current implementation, but Centaur is compatible with other FL algorithms, e.g., FedProx (Li et al., 2020), FedAdam (Reddi et al., 2020), or DM-PFL (Zhang et al., 2023). On the other hand, layer-wise FL methods, such as FedMA (Wang et al., 2020), may have interesting interactions as they also conduct model training in a partitioned fashion.
Second, more algorithmic changes can be done to existing AP training, but we emphasize that whatever we do for AP training can only be useful for applications where all data is available at AP. In many emerging applications, UCDs are the main devices that collect and process data in the first place, and we do not want UCDs to send all the data to AP and put the burden of data selection on the APs shoulders.
Third, we do not perform model selection or architecture search, and our model partitioning simply follows the encoder-classifier style of benchmark networks, but one can enhance the model initialization and partitioning with state-of-the-art advancements in neural architecture search (Li et al., 2021b; Huang et al., 2023).
Fourth, our goal is to present a proof of concept for a multi-device FL that encompasses UCDs and APs. We present a real testbed and perform actual measurements, but in the rest of our experiments, we utilize GPU servers for running end-to-end FL scenarios. The constraints of UCDs are simulated according to the size and type of the model they can train. As a topic for future investigation, and to achieve more precise calculations of compute and storage resources, the end-to-end process of Centaur could be implemented on real-world UCDs and APs.
Fifth, we considered a one-to-one relationship between UCDs and APs, but one can expand Centaur into a many-to-one setting. This, however, needs tackling the challenge of aggregating diverse classifiers trained by various sensing devices.
Sixth, we evaluated Centaur on supervised FL tasks, but it is possible to suggest a supplementary approach for implementing semi-supervised or unsupervised tasks. Loss-based data selection has been studied very well in the ML community and shows promising advantages. Finally, dealing with outliers and distribution shifts are important challenges but we believe these are topics outside the scope of this paper.
9. Conclusions
To include constrained devices in an FL task, we offer an efficient training approach that utilizes data selection to enhance learning while minimizing computational costs. Additionally, we employ partition-based training and aggregation, wherein constrained devices train a specific part of the model, aided by more capable devices that train the complete model. Our assessments validate the benefits of our proposed FL framework, Centaur, in enhancing model accuracy while simultaneously reducing system overhead. Furthermore, we propose several promising directions for future studies in the area of efficient FL on constrained sensory devices.

### v6

@@
 Abstract.
-Federated learning (FL) in multidevice environments creates new opportunities to learn from a vast and diverse amount of private data.
-Although personal devices capture valuable data, their memory, computing, connectivity, and battery resources are often limited.
-Since deep neural networks (DNNs) are the typical machine learning models employed in FL, there are demands for integrating ubiquitous constrained devices into the training process of DNNs.
-In this paper, we develop an FL framework to incorporate on-device data selection on such constrained devices, which allows partition-based training of a DNN through collaboration between constrained devices and resourceful devices of the same client.
-Evaluations on five benchmark DNNs and six benchmark datasets across different modalities show that, on average, our framework achieves $\sim$19higher accuracy and $\sim$58lower latency; compared to the baseline FL without our implemented strategies.
-We demonstrate the effectiveness of our FL framework when dealing with imbalanced data, client participation heterogeneity, and various mobility patterns.
-As a benchmark for the community, our code is available at github.com/dr-bell/data-centric-federated-learning.
-copyright: none
+UbiquitousTo be presented in the 10th ACM/IEEE Symposium on Edge Computing (SEC’25), December 2025, Arlington, VA, USA.
+wearable and mobile devices provide access to a diverse set of data.
+However, the mobility demand for our devices naturally imposes constraints on their computational and communication capabilities.
+A solution is to locally learn knowledge from data captured by ubiquitous devices, rather than to store and transmit the data in its original form.
+In this paper, we develop a federated learning framework, called Centaur, to incorporate on-device data selection at the edge, which allows partition-based training of a deep neural nets through collaboration between constrained and resourceful devices within the multidevice ecosystem of the same user.
+We benchmark on five neural net architecture and six datasets that include image data and wearable sensor time series.
+On average, Centaur achieves $\sim$19% higher classification accuracy and $\sim$58% lower federated training latency, compared to the baseline.
+We also evaluate Centaur when dealing with imbalanced non-iid data, client participation heterogeneity, and different mobility patterns.
+To encourage further research in this area, we release our code at github.com/nokia-bell-labs/data-centric-federated-learning.
+Key words and phrases:
+Federated Learning, Constrained Devices, Data Selection, Partition-Based Training, Edge Intelligence, On-device AI
 1. Introduction
-Multidevice FL enables learning new insights from various clients by training DNNs on personal data in a more private manner (McMahan et al., 2017a; Kairouz et al., 2021).
-However, some wearable or embedded devices have limited computing resources, little on-device storage, and variable network connectivity.
-Most existing FL frameworks do not account for constraints of personal devices and instead assume either modern smartphones (Hard et al., 2018; Niu et al., 2020) or accelerators (Cho et al., 2022; Zhang et al., 2022) as target client devices.
-Although smartphones and accelerators are less powerful than cloud machines, they are nevertheless equipped with mobile GPUs, possess gigabytes of runtime memory, and have fairly stable connectivity to a central server, which simplifies the requirements of DNN training (Yang et al., 2021; Lin et al., 2020).
-Challenge.
-We focus on multidevice FL for ubiquitous constrained devices (UCDs).
-Examples are wearable devices such as earbuds, glasses, or rings, and embedded devices such as environmental or industrial sensors.
-We are driven by the motivation that UCDs are becoming the primary data-producing devices for individuals and industries due to various on-device sensors, such as cameras and inertial measurement units.
-UCDs offer a wealth of spatio-temporal data that is frequently absent when focusing solely on smartphones or plugged-in devices.
-For instance, people often opt for wristbands, smart rings, or earbuds over carrying a smartphone in their hand or pocket while engaging in outdoor, industrial, or sports activities (Seneviratne et al., 2017; Mo and Zhou, 2021).
-Moreover, low-cost sensors distributed across farms or forests remotely monitor crops and livestock and detect potential threats (Papst et al., 2021).
-Such abundance and diversity in sensory data captured by UCDs can facilitate the learning of more effective models for emerging applications in personal and industrial environments.
-Nevertheless, there are two challenges for FL on UCDs:
-The first challenge is the network connectivity to a central server.
-Usually, UCDs use an access point (AP) over wireless channels, e.g., a smartphone or a router, as a relay to connect to the central server.
-Such wireless connections, e.g., via Bluetooth Low Energy (BLE), are unreliable due to the devices’ mobility.
-Particularly for outdoor settings, UCDs may not always be connected to the companion AP, and hence therefore no immediate connection to the central server.
-The second challenge emerges from constraints on the memory and compute resources available on UCDs, particularly the model size and functionality of DNNs that can be trained locally on UCDs.
-Although a naive solution is to transfer all data from UCDs to their respective APs and perform training of DNNs on APs, storing all data could exceed the storage capacity of UCDs, especially when UCDs disconnect from their AP often.
-To tackle these challenges effectively, we require FL frameworks that dynamically distribute data and computations between UCDs and APs.
-Solution.
-We propose an end-to-end FL framework that orchestrates the local training of DNNs among UCDs and APs by integrating on-device data selection alongside partition-based model training.
-We refer to our proposed framework as Centaur.
-(A) Considering computation constraints, Centaur initializes the DNN in two partitions: an encoder (i.e., feature extractor) to be only trained on APs, followed by a lightweight classifier to be trained on both APs and UCDs (§3.2).
-This is inspired by split (Gupta and Raskar, 2018; Thapa et al., 2022) and transfer (Weiss et al., 2016) learning, where UCDs keep the encoder frozen and only train the classifier, while the AP train the full DNN.
-(B) Considering memory and connectivity constraints, Centaur performs data selection (Jiang et al., 2019) by analyzing the training loss and the gradients norm of the classifier part, to decide which data points captured by UCDs contribute more to the training of which part of the DNN.
-Here, data points are categorized as either of (i) discarded if they have very low loss values, (ii) kept locally on the UCD to train the classifier part if their loss values or gradients norm are not very high, and (iii) transmitted to the AP to train both encoder and classifier part if they cause high values for both loss and gradients norm [^1] The assumption is each pair of UCD and AP belongs to the same client, thus transmitting data from a UCD to its AP does not violate the clients’ privacy.
-(§3.3).
-Centaur differs from the prior works on hierarchical FL (Liu et al., 2020; Abad et al., 2020), since they neither perform partition-based training nor incorporate data selection.
-This makes our FL framework adaptive to the available resources in a multidevice setting.
-Evaluation.
-We perform evaluations on five benchmark DNNs and six benchmark datasets.
-We show that compared to existing FL alternatives, Centaur not only saves more bandwidth by reducing the communication cost associated with offloading samples from the UCD to the AP, but also reduces the latency of training, and as a result, the energy consumption, on both the AP and the UCD.
-As a prime example of our experimental results, Figure 1 shows the latency vs.
-accuracy of performing FL for two different tasks: (1) four benchmark DNNs, with their encoder pre-trained on ImageNet (Deng et al., 2009) , and then finally trained for CIFAR10 (Krizhevsky et al., 2009) image classification with 100 clients, and (2) a benchmark ConvNet (Chang et al., 2020), trained from scratch, on three human-activity recognition (HAR) datasets with 20 clients.
+Figure 1.
+The latency of federated training versus classification accuracy on test dataset.
+We compare Centaur with standard federated training that only trains the classifier on ubiquitous constrained devices (UCDs) without data selection and partition-based training.
+(Top plot) Four models were tested on image data from CIFAR-10, and (bottom plot) a model was tested on three datasets of wearable sensor time-series.
+With the growing trend of utilizing ubiquitous devices in personal and industrial environments, federated learning (FL) enables the discovery of new patterns in personal data by training deep neural networks (DNNs) on data captured by various users in a more private manner floriginal; flsurvey.
+Especially, in a multidevice ecosystem, the owner can seamlessly share data across the devices within their trusted environment while protecting data privacy cho2022flame.
+This allows us to utilize the capabilities of multiple devices to participate in the FL process.
+However, wearable and ubiquitous devices have limited computing resources, little on-device storage, and inconsistent network connectivity.
+Most of the existing FL frameworks do not account for constraints of personal devices and instead assume either modern smartphones hard2018federated; niu2020billion, or edge devices with DNN accelerators cho2022flame; zhang2022federated, as target client devices.
+Although smartphones and accelerators are less powerful than cloud machines, they are nevertheless equipped with mobile GPUs, possess gigabytes of runtime memory, and have fairly stable connectivity to a central server, which simplifies the requirements of DNN training yang2021flop; lin2020ensemble.
+In this paper, we propose a novel FL framework to include ubiquitous constrained devices (UCDs), i.e., wearable devices such as earbuds, glasses, or rings, or distributed edge devices such as environmental cameras or industrial sensors.
+We believe UCDs are becoming the primary data-producing devices for both individuals and industries, due to various on-device sensors, such as cameras or inertial measurement units.
+These devices offer a wealth of spatiotemporal data that is frequently absent when focusing solely on smartphones or plugged-in devices.
+For example, people often opt for wristbands, smart rings, or earbuds over carrying a smartphone in their hand or pocket while engaging in outdoor, industrial, or sports activities seneviratne2017survey; mo2021adapting, and sensors distributed across farms or forests remotely monitor crops and livestock to detect hazards papst2021exploring.
+The main motivation of our work is the fact that the abundance and diversity in data captured by ubiquitous devices can facilitate learning more effective models for emerging applications in personal and industrial environments.
+There are several significant challenges.
+Usually, distributed or wearable UCDs use an access point (AP), like a smartphone or a router, as a relay to connect to a central server in the cloud.
+Data captured by a UCD is communicated to the AP using communication protocols like Bluetooth Low Energy (BLE).
+In outdoor settings under mobility, UCDs may not always be connected to the companion AP, thus inaccessible to the central server.
+Beside this, the constrained data storage and computing resources will restrict the size and functionality of DNNs that can be trained locally on UCDs.
+A naive solution is to transfer all data from UCDs to their respective APs and perform training of DNNs on APs.
+Yet, storing all data could exceed the device’s memory, especially when UCDs disconnect from their AP.
+To demonstrate these points, in §3, we conduct a motivating study on a real testbed to identify the essential design requirements that can enable the deployment of FL on resource-constrained devices.
+To tackle these challenges effectively, we believe FL solutions relying on on-device training should dynamically distribute data and computations between UCDs and their companion APs.
+Table 1.
+Comparison with related work.
+(1) Model Partitioning splits the model into an encoder and a classifier.
+(2) Data Selection trains on a subset of available data.
+(3) On-device Training considers backpropagation on constrained devices.
+(4) Spatiotemporal Coverage includes mobility patterns of portable devices.
+Model Partitioning Data Selection On-device Training Spatiotemporal Coverage
+Multitier FL abad2020hierarchical; liu2020client; abdellatif2022communication; cho2022flame; gupta2018distributed × × × ×
+Split FL diao2020heterofl; zhang2021fedzkt; zhang2021parameterized; li2021hermes; dai2022dispfl; bibikar2022federated × × ×
+On-device Training lee2019intermittent; islam2019zygarde; fedorov2019sparse; banbury2021micronets; xu2022etinynet; liberis2021munas; lin2022device; kopparapu2022tinyfedtl × ×
+Centaur (ours)
+In §4 and Figure 3, we introduce Centaur: a federated learning framework that orchestrates local training among UCDs and APs by integrating on-device data selection with partition-based model training.
+(A) Considering computation constraints, we initialize the DNN in two partitions: an encoder (i.e., feature extractor) to be only trained on APs, followed by a lightweight classifier to be trained on both APs and UCDs (§4.1).
+This allows Centaur to utilize all the data available on UCDs for fine-tuning the DNN while occasionally storing and transmitting a portion of the data to the APs to re-train the entire DNN and adjust to the runtime variability in data distributions gupta2018distributed; thapa2022splitfed; weiss2016survey.
+(B) Considering memory and connectivity constraints, we perform data selection jiang2019accelerating by analyzing the training loss and the gradients norm of the classifier part, to decide which data points captured by UCDs contribute more to the training of which part of the DNN.
+Through data selection, data points are categorized as either of (i) discarded if they have very low loss values, (ii) kept locally on the UCD to train the classifier part if their loss values or gradients norm are not very high, and (iii) transmitted to the AP to train both encoder and classifier part if they cause high values for both loss and gradients norm (§4.2).
+We assume that each pair of UCD and AP belongs to the same client, thus transmitting data from a UCD to its AP does not violate the clients’ privacy.
+(C) We benchmark on five DNNs architectures and six datasets of two different modalities, and compared to existing FL alternatives.
+Our results show that Centaur saves more bandwidth by reducing the communication cost associated with offloading samples from a UCD to the AP, and also reduces the latency of training (and accordingly the energy consumption) on both UCD and AP.
+As a prime example of our experimental results, Figure 1 shows the training latency versus classification accuracy of performing FL for two different tasks: (1) four benchmark DNNs, with their encoder pre-trained on ImageNet deng2009imagenet, and then trained for CIFAR10 krizhevsky2009learning image classification with 100 FL clients, and (2) a benchmark ConvNet chang2020systematic, trained from scratch, on three human-activity recognition datasets with 20 FL clients.
 In both figures, we compare Centaur with the baseline of standard FL on UCDs without our implemented strategies.
 Centaur, through data selection and partition-based model training, achieves up to $19\%$ higher accuracy and $58\%$ lower latency, on average.
-We provide thorough experiments and comparisons with other alternative baselines in §6.
+The key contribution of our work is a better integration of ubiquitous constrained devices into federated learning by leveraging the advantages of multi-device ecosystems.
+We achieve this by introducing a customizable data selection scheme and a partition-based training approach, which collectively reduce computational and communication costs while enhancing model accuracy.
+Our experiments demonstrate that Centaur effectively minimizes storage requirements and training time through efficient data selection.
+Furthermore, our analysis accounts for the real-world spatiotemporal mobility of edge devices in FL.
+Empirical evaluations, covering cost, data imbalance, participation heterogeneity, and connection probability, show that Centaur consistently delivers higher efficiency, achieving improved accuracy at lower costs across diverse scenarios.
+2. Related Work
+Table 1 compares Centaur with major prior work on different aspects of multidevice FL.
+Compared to the standard edge–server FL, edge–access point–server FL liu2020client; abad2020hierarchical considers a middle layer, e.g., a cellular base station, to orchestrate FL training across clients and a central server.
+Clients can participate with several devices cho2022flame; diao2020heterofl; zhang2021fedzkt, which achieve better trade-offs between accuracy and bandwidth consumption.
+However, these works assume the local training of the entire model on edge devices, without any resource limitations.
+They also do not utilize the potential of resource-rich devices to assist a resource-constrained device during training.
+They use APs only for aggregation and communication, which makes FL impractical in scenarios involving UCDs.
+Other works consider inference only (i.e., forward pass) fedorov2019sparse; banbury2021micronets; xu2022etinynet; liberis2021munas, sparse training li2021hermes; bibikar2022federated, or heterogenous architectures across clients of different capabilities kang2023nefl; park2024federated.
+Sparse training or pruning requires computational resources (especially memory) that UCDs do not have.
+Few works propose solutions for full on-device training (i.e., forward and backward pass) lin2022device, but not for FL scenarios as they often make impractical assumptions like the availability of large datasets for network architecture search liberis2021munas, or unconstrained memory for training the full model diao2020heterofl; zhang2021fedzkt.
+Transfer learning can be used on UCDs to fix the encoder architecture, which in turn restricts the extraction of newer features, and only trains the last layer for the personalized training zhang2021fedzkt.
+Without utilizing the power of APs, training the entire model on UCDs requires a large sample size to learn generalized features li2021hermes; bibikar2022federated.
+Simply considering different sub-models diao2020heterofl; kang2023nefl; park2024federated requires more sophisticated aggregation strategies and shallower models for resource-constrained devices, which may not be capable of learning the complex patterns from the data.
+Recent works enable training on micro-controllers with limited memory, such as TensorFlow Lite Micro david2021tensorflow or Tiny Training Engine lin2022device.
+These works do not consider FL as a potential use case and do not account for the advantages of a multidevice ecosystem.
+3. A Motivating Study
+As depicted in Figure 3, a ubiquitous constrained device (UCD) is a small, low-power portable device with restricted hardware and connectivity resources.
+Examples include smart rings, earbuds or glasses (personal wearables) and tiny environmental or industrial sensors.
+Each UCD connects to the user’s personal access point (AP), usually over a low-power link such as Bluetooth Low Energy.
+An AP is a resourceful device that has much better hardware and connectivity capabilities than its corresponding UCD (for instance, a smartphone, home router, or gateway).
+We assume the AP is always part of the user’s device ecosystem that allows the UCD to transmit data to the AP without violating the user’s privacy.
+We generally assume one AP serves multiple UCDs (e.g., a person’s phone might pair with a smartwatch, earbuds, and other sensors simultaneously).
+Indeed, modern homes routinely host on the order of 20–30 connected devices, so a single AP often handles many UCDs.
+A server is a central entity owned by a service provider that orchestrates FL over several clients.
+Considering a deep neural network architecture, an encoder ($E$) is the first part of the DNN that extracts the features from input data.
+A classifier ($C$) is the second part of the DNN, which performs the final classification using the encoded features.
+We consider supervised learning tasks where $\mathbf{X}$ denote the input data and $\mathbf{y}$ denote its label.
+Setup.
+To identify the key design choices, we create a testbed with four RaspberryPi 4 Model B as edge devices velasco2018performance; drakopoulos2019real; all running the 64-bit Raspbian OS with total secondary storage of 3232GB.
+We set up four clients with varying primary memories: C1, C2, C3, and C4 having 11GB, 22GB, 44GB, and 22GB, respectively.
+Indeed, Raspberry Pi devices are more capable than many wearables; however, they serve as a widely adopted, reproducible proxy for constrained clients cho2022flame, enabling controlled CPU/memory benchmarking while we contextualize our findings with real-world wearable hardware specifications and multitasking considerations.
+For analyzing the CPU and memory consumption of the clients, we use logs of /proc/stat and free operating-system calls, respectively.
+We use MobileNetV3 howard2019searching, pre-trained on ImageNet, as the encoder followed by two fully-connected layers with an intermediate drop-out layer.
+Using CIFAR10, we measure the resources and time needed for both on-device training (per epoch) and inference.
+We implement FL with Flower beutel2022flower and use a desktop computer as the server, as shown in Figure 2a.
 Figure 2.
-Training latency vs.
-test accuracy of Centaur compared with standard UCD training (i.e., only training the classifier on UCDs without data selection and partition-based training).
-(Left) Four benchmark pre-trained DNNs fine-tuned on CIFAR10 dataset with FL.
-(Right) A ConvNet trained with FL on three benchmark wearable sensor datasets: MotionSense, PAMAP2, and UCIHAR.
-Contribution.
-Our solution enables UCDs to efficiently participate in multidevice FL alongside their upstream APs, by incorporating a customizable data selection scheme and a partition-based training regime.
-The impact is a decrease in computational and communication costs, as well as an improvement in the model’s test accuracy.
-Our experiments on a small testbed of RaspberryPis validate that Centaur is promising in decreasing overall storage requirements and training time due to employing effective data selection.
-Moreover, our analysis on the mobility of UCDs accounts for the real-world spatiotemporal mobility of edge devices participating in FL.
-Our extensive empirical analyses, including measurement of cost, imbalance data, participation heterogeneity, and connection probability, indicate that Centaur achieves higher efficiency (i.e., high accuracy and low cost) across most scenarios.
-Our code is open-source to encourage further advancements in this domain.
-2. Identifying Design Challenges: A Motivational Study
-Terminology.
-A ubiquitous constrained device (UCD) is a mobile device with restricted hardware and connectivity resources.
-An access point (AP) is a resourceful device that has much better hardware and connectivity capabilities than its corresponding UCD.
-A server is a central entity owned by a service provider that organizes FL over several clients.
-An encoder ($E$) is the first part of the DNN and extracts the features of input data.
-A classifier ($C$) is the second part of the DNN and classifies the encoded features into a final prediction.
-Let $\mathbf{X}$ denote the input data and $\mathbf{y}$ denote its label.
-2.1. The Study
-We perform measurement experiments to identify the key design requirements that can facilitate the deployment of FL on constrained devices.
-Setup.
-We create a testbed with four RaspberryPi 4 Model B as the clients (Velasco-Montero et al., 2018; Drakopoulos et al., 2019); all running the 64-bit Raspbian OS with total secondary storage of $32$GB.
-We set up a cohort of clients having varying primary memories, with clients C1, C2, C3, and C4 having $1$GB, $2$GB, $4$GB, and $2$GB of primary memory, respectively.
-For analyzing the CPU and memory consumption of the clients, we use the logs from the /proc/stat and free system calls, respectively.
-For the model architecture, we use a benchmark architecture MobileNetV3 (Howard et al., 2019) (pre-trained on ImageNet) as the encoder followed by a medium-sized classifier with two fully connected layers with an intermediate drop-out layer.
-We use CIFAR10 dataset and measure the resource and time consumed on-device for both training, per each epoch, and inference.
-Finally, for implementing FL setup, we use Flower (Beutel et al., 2022) over four RaspberryPi devices and a Mac Mini as the server, as shown in Figure 2a.
-Running Standard FL.
-We implement standard FL (for $100$ rounds) over four RaspberryPi clients with the standard FedAvg (McMahan et al., 2017b).
-We consider two scenarios: (1) when only the classifier part is trained, and (2) when the entire model (encoder and classifier) is trained.
-For the first scenario, we observed in our experiments that the standard FL runs successfully over all four clients.
-However, for the second scenario, client C1 with 1GB of primary memory fails to participate in the FL process.
-Notably, 1GB of primary memory is still significant in terms of memory capability and typical UCDs usually have a much lesser amount of memory than this.
+On four RaspberryPis, we measure memory consumption of running FL for when (i) only the classifier is trained, (ii) the entire model is trained, and (iii) the entire model runs inference only.
+[…]
+Key Observations.
+We run FedAvg floriginal for $100$ rounds with four RaspberryPi devices, considering two scenarios: (1) when only the classifier part is trained, and (2) when the entire model (encoder and classifier) is trained.
+When running our experiments for the first scenario, we observed that the standard FL runs successfully over all four clients.
+However, for the second scenario, client C1 with 1GB of primary memory always failed to participate in the FL process.
+This is primarily due to the significant rise in the CPU and memory consumption when the entire model (both encoder and classifier) are trained on the device (see Figure 2).
+Notably, 1GB of primary memory is still significant in terms of memory capability, and many UCDs may have significantly less memory than this.
 Therefore, it is clear that in such cases, it is much more challenging to deploy standard FL over UCDs.
-To further assess this, we conduct a series of experiments in the following section to investigate the resource and time consumption for training DNNs on our testbed.
-On-device Training.
-Considering the challenges discussed previously, we analyze the latency and resource consumption for local training of MobileNetV3.
-Figure 2 shows that there is a significant rise in the overall CPU and memory consumption when the entire model (encoder and classifier) is trained on-device in comparison to when only the classifier is trained and the encoder kept frozen.
-A key observation is that by training only the classifier, we require less processing power and approximately the same amount of memory resources as when the entire model is used solely for inference.
-Moreover, Figure 3 shows that the time taken to train the classifier is less than the time to perform a single forward pass on the entire model.
-These measurements confirm that training the classifier only is 50×50$\times$50 × time efficient compared to training the entire model, which helps in reducing the battery consumption significantly.
-When we follow a naive approach of randomly dropping samples from the training dataset, we observe a decline in the overall training time of the classifier (see Figure LABEL:fig:select_train).
-However, this comes at the cost of reduced accuracy, which might be detrimental to the overall system.
-Figure 2.
-Resource consumption of training DNNs on (a) the designed setup over four RaspberryPi 4 Model B.
-We compare (b) CPU Consumption and (c) Memory Consumption for three scenarios when: (i) only the classifier is updated, (ii) the whole model is updated, and (ii) the whole model is used only at inference mode.
-2.2. Insights
-Here, we summarize the understandings derived from our motivational study:
-(1) Resource and Training Efficiency.
-As UCDs typically have limited memory, training on the device should be memory efficient.
-This motivates using a frozen pre-trained encoder, and only training the classifier part on the device (Mathur et al., 2021).
-However, freezing the weights of the encoder limits the expressiveness of the model and prevents the model from learning important features of the newly collected data over the UCDs.
-Thus, while minimizing the on-device model training on the UCDs, our system should ensure that the encoder is periodically trained on a portion of new data collected from the UCDs, potentially on a more resourceful device where the entire model can be trained.
-(2) Data Efficiency and Spatiotemporal Coverage.
-Typically, UCDs are deployed to maximize the spatiotemporal coverage of activities and events.
-Hence, our design should leverage the spatiotemporal richness of such data to train a better model.
-Such spatiotemporal diversity also makes UCDs more prone to intermittent connectivity, which may require the framework to store training data on-device.
-UCDs possess limited secondary storage, and even when they have persistent connectivity to the AP, streaming high volumes of data could drain their battery.
-Thus, our design should dynamically select a small subset of data points for storing locally or transmitting to the APs.
-This needs to be designed in a sophisticated manner such that it does not compromise the overall performance of the system.
-Figure 3.
-(a) Time for backpropagation across (i) the entire model (encoder and classifier), (ii) across the classifier only, and (iii) for forward pass on the entire model.
-(b) The trend of training time and accuracy (for one epoch) for different sizes of training sets.
-3. Our Proposed Framework: Centaur
-With our understanding developed from the motivational study, we introduce our proposed framework. Figure 4 gives a schematic overview, and below we elaborate on the algorithms used in Centaur.
-3.1. The Overview
-The first level includes UCDs, which are connected to their APs via BLE, Wi-Fi, Backscatter, or other wireless technologies.
-We assume the connection between UCD and AP is not always available or reliable due to the natural mobility of UCDs.
-We consider that the data selection procedure and the training of the classifier only happen on UCDs.
-The second level, which includes APs, is responsible for training the encoder in conjunction with the classifier atop it.
-The third level includes a central server, which performs model aggregation for both classifiers and encoders that are trained by UCDs and APs.
-Here, we present a step-by-step description of our proposed framework.
-Step \⃝raisebox{-0.9pt}{1}: We assume that the server receives the information about UCDs’ resource.
-Accordingly, the server initializes a DNN model $M$ that is formed of an encoder $E$ and a classifier $C$, and then distributes the $E$ and $C$ to the APs and subsequently to UCDs.
-We explain model initialization details in §3.2.
-Step \⃝raisebox{-0.9pt}{2}: UCDs start FL by performing data selection on the collected data $D$, based on the computed loss value via the forward pass and the magnitude of the gradient of the classifier’s last layer.
-Each sample point will be categorized in one of (i) $D_{N}$ to be discarded, (ii) $D_{C}$ for locally training $C$ on the UCD, or (iii) $D_{M}$ to be transmitted to AP for training $M$.
-We explain data selection details in §3.3.
-Step \⃝raisebox{-0.9pt}{3}: The UCD discards data $D_{N}$ in the current epoch without any further training steps or backward pass on them.
-Step \⃝raisebox{-0.9pt}{4}: The UCD performs backward pass through the classifier $C$ on data $D_{C}$, while the encoder $E$ is frozen.
-Then, the UCD updates the parameters of $C$ via computed gradients.
-Notice that, in practice, the gradient computation done in Step \⃝raisebox{-0.9pt}{2} can be reused for this update.
-Step \⃝raisebox{-0.9pt}{5}: UCDs share their updated classifier $C$ to the server, and the server performs aggregation on all $C$ from participating UCDs to obtain one aggregated classifier.
-Step \⃝raisebox{-0.9pt}{6}: The server sends aggregated classifier back to the UCDs and APs.
-Step \⃝raisebox{-0.9pt}{7}: With the updated classifier, the AP trains the encoder $E$ together with the classifier $C$ atop $E$ on data $D_{M}$.
-That is, $D_{M}$ is fed for a forward pass and then a backward pass throughout to update $M$.
-Step \⃝raisebox{-0.9pt}{8}: The updated $M$ (including $E$ and $C$) is shared from the AP to the server.
-After collecting updated $M$ from all participating APs, the server performs model aggregation to obtain the aggregated encoder $E$ and classifier $C$ and distribute them back to UCDs and APs as in Step \⃝raisebox{-0.9pt}{6}.
-See § 3.4 for more details of FL training procedures.
-Steps from \⃝raisebox{-0.9pt}{2} to \⃝raisebox{-0.9pt}{8} will be repeated for several FL rounds to update $C$ on UCDs and $E$ on APs.
-Figure 4.
-The overview of our framework, including model initialization (§ 3.2), data selection (§ 3.3), and partition-based training and aggregation (§ 3.4).
-We explain the set up in § 3.1 and elaborate the details of Steps \⃝raisebox{-0.9pt}{1} to \⃝raisebox{-0.9pt}{8}.
+Moreover, training only the classifier consumes far less resources and is even comparable to the resource usage required for a forward pass.
+We noticed a similar pattern for time consumption, where training the classifier only is $50\times$ time efficient compared to training the entire model.
+One alternative solution to decrease the overall training time is to reduce the training data size by randomly discarding samples.
+However, we observe that such an approach comes at a cost of significantly reduced accuracy, which is detrimental to the overall system.
+4. Method
+In Figure 3, we introduce Centaur, our federated learning framework. Here, we elaborate on each component in detail.
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Since full backpropagation involves a forward pass, the storage of all intermediate activations, and a backward pass through every layer to compute gradients, it is the most compute- and memory-intensive phase of model training.
+Centaur mitigates this by freezing the deep encoder and training only the lightweight classifier head on UCDs.
+Each device performs a single forward pass through the fixed encoder and backpropagates only through the shallow fully-connected classifier, which has orders of magnitude fewer parameters.
+This design significantly reduces the on-device workload.
+Our motivationa study in §3 confirmed this trend, with classifier-only training running several times faster and using far less memory than updating the entire network.
+Thus, Centaur enables effective training on local data while keeping UCD training lightweight and efficient.
 4.1. Model Initialization
-To partition a DNN into an encoder $E$ and a classifier $C$, one option is to use pre-trained benchmark DNNs; considering numerous well-trained models, e.g., in Model Zoo (Koh, 2019).
-Usually, DNN architectures from a candidate space of pre-trained models have already been well developed, with efficient quantization and compression capabilities, to be deployed on UCDs for inference (Fedorov et al., 2019; Banbury et al., 2021; Xu et al., 2022; Liberis et al., 2021; Malekzadeh et al., 2021; Li et al., 2021a).
+To partition a DNN into an encoder $E$ and a classifier $C$, one option is to use pre-trained benchmark DNNs; considering numerous well-trained models.
+Usually, DNN architectures from a candidate space of pre-trained models have already been well developed, with efficient quantization and compression capabilities, to be deployed on UCDs for inference fedorov2019sparse; banbury2021micronets; xu2022etinynet; liberis2021munas; li2021hermes.
 We aim to utilize the full capability of UCDs for on-device training of the classifier $C$.
@@
 To achieve these, we can iteratively look into candidate architectures that fit in the limited memory available on these UCDs.
-A robust mechanism in this direction is to obtain a history of memory usage of a particular device which reflects the typical memory availability on the UCD.
+A robust mechanism in this direction is to obtain a history of memory usage of a particular device, which reflects the typical memory availability on the UCD.
+To make the selection process explicit, we consider a set of lightweight architectures $\{m_{k}\}$: a small set of widely used mobile convolutional models (e.g., SqueezeNet and MobileNet-style variants), alongside fully connected networks with varying depth (1–3 layers) and width (16–128 neurons).
+These architectures are chosen because they span the typical memory–compute trade-off space encountered in edge devices while still supporting on-device gradient updates.
+Among the feasible candidates that satisfy the device’s resource constraints, we select the smallest model that achieves an accuracy within a small tolerance of the best-performing candidate.
+This strategy effectively balances efficiency and performance, ensuring that the chosen classifier maximizes on-device feasibility without incurring a significant loss in learning effectiveness.
 Formally speaking, the iterative process can be viewed as an optimization problem described as
-(1) $C=\min\{m_{1},m_{2},\ldots,m_{k}|\mathbb{P}(m_{k})\times\mathbb{B}<\mathbb{M}\},$
-where $\mathbb{P}(m_{k})$ provides the number of parameters of the candidate architecture $m_{k}$, $\mathbb{B}$ is the number of bytes used to store the parameters and intermediate results in computations, and $\mathbb{M}$ is the memory available in the device for FL participation.
-The parameter $\mathbb{M}$ can be measured and logged as empirical observations of the UCD memory usage, to be sent to the server as we discussed in Step \⃝raisebox{-0.9pt}{1} in Figure 4.
-In this paper, we limit our classifier architectures with fully connected layers; thus, we find a solution by either reducing the number of neurons or the layers to find a potential architecture for $C$ that can fit into the requirements for $\mathbb{M}$ in Equation (1).
-In the same vein, we tackle the problem of low computing resources by choosing the smallest architecture for $C$ and then deploying it on the UCDs.
+$C=\operatorname*{argmin}_{\{m_{1},\dots,m_{K}\}}\;\mathbb{P}(m_{k})\,\mathbb{B}\penalty 10000\ \penalty 10000\ \text{s.t.}\penalty 10000\ \penalty 10000\ \mathbb{P}(m_{k})\,\mathbb{B}\leq\mathbb{M}\penalty 10000\ \text{and}\penalty 10000\ \text{acc}(m_{k})\geq\tau^{*},$ (1)
+[…]
+21Inputs: (1) $\mathcal{A}$: the set of AA clients each having two devices, one UCD and one AP, (2) $D^{k}$: local dataset for each client $a\in\{1,\dots,A\}$, (2) $f_{E}$: the encoder part with parameters $W_{E}$, (3) $f_{C}$: the classifier part with parameters $W_{C}$, (4) $R$: the total number of FL rounds.
+30.1cm
+54 for $r\in\{1,\dots,\nicefrac{{R}}{{2}}\}$ rounds do
+7 6 $K$ clients $\Leftarrow$ randomly select $K$ clients from a total of AA, using uniform sampling
+8 0.1cm
+9  — Training on UCDs —
+11 10 for $k\in\{1,...,K\}$ do
+13    12 $\mathbf{X}^{k},\mathbf{y}^{k}\leftarrow D^{k}$ (all the samples in the local dataset of client $k$ UCD)
+15    14 $\ell^{k}=\mathcal{L}(f_{E}\circ f_{C}(\mathbf{X}^{k}),\mathbf{y}^{k})$ (forward pass to compute per-sample loss values)
+17    16 $D_{N}^{k},D_{C}^{k},D_{M}^{k}\Leftarrow$ based on $\ell^{k}$, perform data selection as detailed in Section 4.2
+19    18 for all $(\mathbf{X}^{k}_{j},\mathbf{y}^{k}_{j})\in D_{C}^{k}$ do
+21       20 $g^{k}_{C}={\partial\ell^{k}}/{\partial W_{C}}$ (compute $f_{C}$’s gradients on $D_{C}^{k}$ to update $f_{C}$)
+23       22 $W_{C}^{k}=Optimizer(W_{C}^{k},g^{k}_{C})$ (update $f_{C}$’s parameters using gradients)
+25       24 $D^{+}\Leftarrow$ based on $g^{k}_{C}$, perform data selection as detailed in Section 4.2
+27       26 $D_{M}^{k}=D_{M}^{k}\cup D^{+}$ (enhance $D_{M}^{k}$ by high-value gradients data points)
+28  — Server Side Aggregation on the Classifier Part —
+30 29 $W_{C}\leftarrow\frac{1}{K}\sum_{k\in K}W_{C}^{k}$
+32 31 $W_{C}^{k}\leftarrow W_{C}$ (update all clients with new $W_{C}$)
+33 0.1cm
+34  — Training on APs —
+36 35 for $k\in\{1,\dots,K\}$ do
+37
+39    383pt $\mathbf{X}^{k},\mathbf{y}^{k}\leftarrow D_{M}^{k}$ (all selected samples for training $f_{E}$ on client $k$ AP)
+41    40 $\ell^{k}=\mathcal{L}(f_{E}\circ f_{C}(\mathbf{X}^{k}),\mathbf{y}^{k})$ (forward pass to compute per-sample loss values)
+43    42 $g^{k}={\partial\ell^{k}}/{\partial(W_{E},W_{C})}$ (compute gradients on $D_{M}^{k}$ to update $f_{E}$ and $f_{C}$)
+45    44 $W_{E}^{k},W_{C}^{k}=Optimizer(W_{E}^{k},W_{C}^{k},g^{k})$ (update all parameters using gradients)
+46  — Server Side Aggregation on both Encoder and Classifier Parts —
+48 47 $\{W_{E},W_{C}\}\leftarrow\frac{1}{K}\sum_{k\in K}\{W_{E}^{k},W_{C}^{k}\}$
+50 49 $\{W_{E}^{k},W_{C}^{k}\}\leftarrow\{W_{E},W_{C}\}$ (update all clients with the new aggregated model)
+51 return $\{W_{E},W_{C}\}$
+Algorithm 1 Centaur: Multidevice Federated Learning via Partition-based Training and Data Selection
 4.2. Data Selection
@@
 4.2.1. Loss-based Selection.
-All available data $D$ at round $r$ are fed into the model $M$ to compute their loss values $\ell_{r}(f_{M}(\mathbf{X}),\mathbf{y})$.
-We consider lower loss values as an indication of being less important, and higher loss values show the importance of the data (Jiang et al., 2019).
 […]
-(2) $\begin{cases}\mathcal{P}_{N}(\ell_{r+1})=1-[\text{CDF}^{\ell}_{r}(\ell_{r+1})]^{\alpha}\\ \mathcal{P}_{M}(\ell_{r+1})=[\text{CDF}^{\ell}_{r}(\ell_{r+1})]^{\beta}\end{cases}$
-where $\alpha$ and $\beta$ parameters determine the level of selectivity.
-At round $r+1$, for a sample with loss value $\ell_{r+1}$, the sample is selected into $D_{N}$ with probability of $\mathcal{P}_{N}(\ell_{r+1})$.
-This implies that the lower the loss value for a sample, the greater the probability it will be discarded.
-Similarly, $\mathcal{P}_{M}$ is used to select samples with the highest losses as $D_{M}$; which means that the higher the loss value for a sample, the greater the probability it will be chosen for training the entire model.
+$\begin{cases}\mathcal{P}_{N}(\ell_{r+1})=1-[\text{CDF}^{\ell}_{r}(\ell_{r+1})]^{\alpha}\\ \mathcal{P}_{M}(\ell_{r+1})=[\text{CDF}^{\ell}_{r}(\ell_{r+1})]^{\beta}\end{cases}$ (2)
 […]
 4.2.2. Gradient-based Selection.
-[…]
-(3) $\mathcal{P}_{M}^{+}(\|g_{r+1}\|)=[\text{CDF}^{g}_{r}(\|g_{r+1}\|)]^{\gamma}$
+For samples that are not selected through $\mathcal{P}_{N}$ and $\mathcal{P}_{M}$, and to avoid consuming extra resources of UCDs, we can further compute the gradients of the classifier’s last layer.
+The norm of these gradients gives us a useful hint about the sample’s importance while requiring much less computation than computing all layers’ gradients katharopoulos2018not.
+The last layer’s norm at round $r$ is produced during training of the classifier by $g_{r}=\frac{\partial\ell_{r}}{\partial W_{r}}$, where $W_{r}$ is the weights of the classifier $C$’s last layer at round $r$.
+Following the same idea, we derive a CDFg for gradient norm values and build a queue to keep track of the computed norm of gradient values.
+Samples with larger gradients have larger impacts on the model’s weights; thus, we also keep these samples for training the entire model $M$ as $E$ might learn new “features” from them.
+Thus, the probability of adding a sample to $D_{M}$ at round $r+1$ is defined as:
+$\mathcal{P}_{M}^{+}(\|g_{r+1}\|)=[\text{CDF}^{g}_{r}(\|g_{r+1}\|)]^{\gamma}$ (3)
 where $\gamma$ is used for customizing the selection rate.
@@
 Moreover, one may substitute this module with a different data selection technique that could also be dependent upon the specific use case.
-321 Inputs: (1) $\mathcal{A}$: the set of $A$ clients each having two devices, one UCD and one AP, (2) $D^{k}$: local dataset for each client $a\in\{1,\dots,A\}$, (2) $f_{E}$: the encoder part with parameters $W_{E}$, (3) $f_{C}$: the classifier part with parameters $W_{C}$, (4) $R$: the total number of FL rounds.
-for
-$r\in\{1,\dots,\nicefrac{{R}}{{2}}\}$ rounds
- do
-6      54 $K$ clients $\Leftarrow$ randomly select $K$ clients from a total of $A$, using uniform sampling — Training on UCDs — for $k\in\{1,...,K\}$ do
-11            10987 $\mathbf{X}^{k},\mathbf{y}^{k}\leftarrow D^{k}$ (all the samples in the local dataset of client $k$ UCD) $\ell^{k}=\mathcal{L}(f_{E}\circ f_{C}(\mathbf{X}^{k}),\mathbf{y}^{k})$ (forward pass to compute per-sample loss values) $D_{N}^{k},D_{C}^{k},D_{M}^{k}\Leftarrow$ based on $\ell^{k}$, perform data selection as detailed in Section 3.3 for all $(\mathbf{X}^{k}_{j},\mathbf{y}^{k}_{j})\in D_{C}^{k}$ do
-16                  15141312 $g^{k}_{C}={\partial\ell^{k}}/{\partial W_{C}}$ (compute $f_{C}$’s gradients on $D_{C}^{k}$ to update $f_{C}$)$W_{C}^{k}=Optimizer(W_{C}^{k},g^{k}_{C})$ (update $f_{C}$’s parameters using gradients) $D^{+}\Leftarrow$ based on $g^{k}_{C}$, perform data selection as detailed in Section 3.3 $D_{M}^{k}=D_{M}^{k}\cup D^{+}$ (enhance $D_{M}^{k}$ by high-value gradients data points)
-20      191817 — Server Side Aggregation on the Classifier Part — $W_{C}\leftarrow\frac{1}{K}\sum_{k\in K}W_{C}^{k}$ $W_{C}^{k}\leftarrow W_{C}$ (update all clients with new $W_{C}$) — Training on APs — for $k\in\{1,\dots,K\}$ do
-25            24232221 $\mathbf{X}^{k},\mathbf{y}^{k}\leftarrow D_{M}^{k}$ (all selected samples for training $f_{E}$ on client $k$ AP) $\ell^{k}=\mathcal{L}(f_{E}\circ f_{C}(\mathbf{X}^{k}),\mathbf{y}^{k})$ (forward pass to compute per-sample loss values) $g^{k}={\partial\ell^{k}}/{\partial(W_{E},W_{C})}$ (compute gradients on $D_{M}^{k}$ to update $f_{E}$ and $f_{C}$) $W_{E}^{k},W_{C}^{k}=Optimizer(W_{E}^{k},W_{C}^{k},g^{k})$ (update all parameters using gradients)
-28      2726 — Server Side Aggregation on both Encoder and Classifier Parts — $\{W_{E},W_{C}\}\leftarrow\frac{1}{K}\sum_{k\in K}\{W_{E}^{k},W_{C}^{k}\}$$\{W_{E}^{k},W_{C}^{k}\}\leftarrow\{W_{E},W_{C}\}$ (update all clients with the new aggregated model)
-29 return $\{W_{E},W_{C}\}$
-Algorithm 1 Multidevice FL with Data Selection and Partition-based Training
+While our proposed data selection strategy is primarily empirical, it is motivated by the theoretical principles of importance sampling in deep learning.
+Prior work has shown that sampling data points in proportion to their loss values or gradient magnitudes can approximate the optimal sampling distribution that minimizes the variance of stochastic gradient estimates and thus accelerates convergence katharopoulos2018not; jiang2019accelerating.
+In particular, high-loss samples tend to carry larger gradient norms and contribute more substantially to the update direction, while samples with small losses or small gradient norms contribute less.
+Our combination of loss value and last-layer gradient norm therefore aligns with these theoretical insights: it adaptively emphasizes samples expected to produce high-informative gradients, balancing computational cost with learning efficacy.
+Nonetheless, we acknowledge that our current formulation does not provide a formal convergence guarantee.
+Developing theoretical guarantees through variance-reduction or convergence-rate analysis is a promising direction for future work.
 4.3. Partition-based Training and Aggregation
-The training process is conducted on both UCDs and APs.
+The results of §3 highlight why we freeze the encoder and only train the classifier on UCDs.
+Training just the classifier consumed far less memory and CPU, comparable to running an inference pass, and ran much faster than full-model training.
+Since most UCDs have far less memory than even a RaspberryPi (e.g.
+a few MB of storage or only a few hundred KB of SRAM), full-model training on UCDs is generally infeasible.
+By freezing the encoder on UCDs, Centaur aims to reduce the on-device computation and memory requirements, allowing each UCD to train only the compact classifier locally.
+The more powerful AP (smartphone or router) can then periodically update the encoder on the offloaded samples, so that the overall model can still learn richer features even though each UCD only fine-tunes the classifier part.
+Therefore, our training is conducted on both UCDs and APs.
 While UCDs only train the classifier $C$ on data $D_{C}$, APs train the complete model $M$ on data $D_{M}$.
-Such procedure allows the model to be (partially) updated when UCDs are offline, in addition to the full updates when UCDs have a connection to the internet.
+Such a procedure allows the model to be (partially) updated when UCDs are offline, in addition to the full updates when UCDs have a connection to the internet.
 This better utilizes UCDs’ spatiotemporal richness and consequently improves the training efficiency in our design requirements.
-Algorithm 1 shows the process of partition-based training and aggregation in Centaur.
-We set the algorithm to run $\nicefrac{{R}}{{2}}$ rounds, considering that UCD training and AP training proceed iteratively in one round.
+Algorithm 1 shows partition-based training and aggregation in Centaur, running $\nicefrac{{R}}{{2}}$ rounds, considering that UCD training and AP training proceed iteratively in one round.
 All training is conducted on the client’s devices either on UCDs or APs.
-Then with all updated weights, the server performs Federated Averaging (FedAvg) (McMahan et al., 2017b) to obtain one aggregated global model.
+Then with all updated weights, the server performs Federated Averaging (FedAvg) floriginal to obtain one aggregated global model.
 The new global model needs to be distributed to all clients UCDs and APs such that (i) the training on AP has an updated classifier, and (ii) the training on UCD in the next round has both an updated encoder and classifier.
-4. Metrics
-To evaluate the performance of our FL framework, compared to other FL alternatives, we use the following metrics:
-1) Accuracy.
-This is the classification accuracy of the trained model on a test set hosted by the central server.
-For a model obtained at the end of each training round, the test accuracy is computed as
-$acc\equiv\frac{1}{T}\sum_{i=1}^{T}\mathbb{1}\big{(}\operatorname*{argmax}\big{(}f_{E}\circ f_{C}(\mathbf{X}_{i})\big{)}=\mathbf{y}_{i}\big{)},$
-where test set has $T$ pairs of $(\mathbf{X}_{i},\mathbf{y}_{i})$ and $\mathbb{1}(\mathrm{C})$ denote the indicator function that outputs 1 if condition $\mathrm{C}$ holds.
-2) Multiply–Accumulate (MAC).
-This operation is a common step that computes the product of two numbers and adds that product to an accumulator ($a\leftarrow a+(b\times c)$); a fundamental operation for both any DNN layers during training and inference.
-We use fvcore (FAIR, 2019) library to compute the number of MAC operations.
-Since fvcore only supports counting MACs in a forward pass, and to count MACs in a backward pass, we use the heuristic that FLOPs (i.e., double MACs) ratio of the backward-forward pass is typically between 1$\times$ and 3$\times$ and most often is 2$\times$ based on models’ specific layer types, according to previous observations (Marius Hobbhahn, 2021).
-3) Bandwidth.
-As the model size (i.e., the encoder or the classifier) can be different in each FL round, we use fvcore to count the number of parameters that are communicated in each round.
-Based on the model size and number of communications among UCDs, APs, and the server, we compute the amount of bandwidth that is consumed.
-When it is a case, we also count the number of sample points that are uploaded from UCDs to APs.
-4) Latency.
-The latency usually has a linear relation with MAC operations; due to the lack of specialized accelerators (Liberis et al., 2021; Lin et al., 2022).
-Thus, we estimate the latency of model training based on the processor’s frequency:
-(4) $\text{Latency}=c*\frac{\text{Total MACs}}{\text{Processor Frequency}}.$
-The ratio between MAC operations and the processor’s instructions, $c$, is typically between $1$ and $2$ based on specific instruction sets/compilers.
-For simplicity, we assume that each MAC operation translates to two instructions in an MCU (i.e., $c=2$).
-Notice that existing processors (e.g., Intel’s Load Effective Address) completes one MAC in one instruction (WiKiBooks, 2021)), and this only scales the experimental results and does not change the final conclusions.
-Finally, the latency of communication is estimated based on the total amount of data needed to be transmitted, divided by the up-link speed or the down-link speed of devices.
-5) Energy.
-This is the total execution time multiplied by the processor’s consumed power per unit of time.
-The energy consumption of communication can also be calculated based on the total time of transmitting data multiplied by the transmitters’ power per unit of time.
-We set the values for UCDs and APs in our simulations as follows:
-(A) As computation resources, we set up:
-Device
-CPU Frequency (MHz)
-Storage Capacity (MB)
-Power Capacity (mW/MHz)
-Timeout (s)
-UCD 100 5 0.05 Inf
-AP 2000 4096 1.5 Inf
-(B) For communication resources, we set up:
-Device
-Uplink(Mbit/s)
-Downlink (Mbit/s)
-Communication Energy (W)
-Disconnect Probability
-UCD 2 2 0.0001 0.5
-AP 10 100 10 0
+Centaur is the first FL framework to simultaneously combine model partitioning, data selection, and on-device training.
+Our approach of partitioning the DNN into encoder and classifier, selecting a subset of data on-device, and performing lightweight on-device updates, has not been addressed jointly in earlier work.
+In contrast, existing hierarchical FL schemes either train full models on each device or only aggregate updates at the AP without splitting the model.
 5. Experimental Setup
@@
 We use six commonly used datasets.
-CIFAR10 and CIFAR100 (Krizhevsky et al., 2009) that both include 50K training samples and 10K test samples, with 10 and 100 classes respectively.
-EMNIST (Cohen et al., 2017) that includes 112K training samples with 47 classes.
-UCIHAR (Anguita et al., 2013) is a widely used dataset of 30 users performing 6 daily activities; data from the accelerometer and gyroscope sensors were collected by a smartphone worn on the waist.
-Data from 21 users are used for training and that of the other 9 users for testing purposes.
-MotionSense (Malekzadeh et al., 2019) also includes accelerometer and gyroscope data from 24 users with a smartphone in the pocket of the trousers who performed 6 activities in 15 trials.
+CIFAR10 and CIFAR100 krizhevsky2009learning that both include 50K training samples and 10K test samples, with 10 and 100 classes, respectively.
+EMNIST cohen2017emnist that includes 112K training samples with 47 classes.
+UCIHAR Anguita2013APD is a widely used dataset of 30 users performing 6 daily activities; data from the accelerometer and gyroscope sensors were collected by a smartphone worn on the waist.
+Data from 21 users is used for training, and that of the other 9 users for testing purposes.
+MotionSense malekzadeh2019mobile also includes accelerometer and gyroscope data from 24 users with a smartphone in the pocket of the trousers who performed 6 activities in 15 trials.
 We use as test data one trial session for each user and as training data the remaining trial sessions (e.g., one trial of “walking” of each user is used as test data and the other two trials are used as training).
-PAMAP2 (Reiss and Stricker, 2012) dataset contains data of 13 different physical activities, performed by 9 subjects wearing 3 devices and a heart rate monitor.
-We use the training-test split provided by PersonalizedFL library (Lu and Wang, 2023).
+PAMAP2 reiss2012introducing dataset contains data of 13 different physical activities, performed by 9 subjects wearing 3 devices and a heart rate monitor.
+We use the training-test split provided by the PersonalizedFL library PersonalizedFL.
 Models.
-For image classification datasets, we select four DNNs commonly used in mobile/edge-oriented literature: (1) EfficientNet-v2 (Tan and Le, 2021), (2) MobileNet-v3 (Howard et al., 2019), (3) MNASNet (Tan et al., 2019), and (4) ShuffleNet (Ma et al., 2018), with 5.3M, 2.2M, 2.5M, and 1.4M number of parameters respectively.
-For human-activity recognition datasets, we borrow the ConvNet architecture proposed in (Chang et al., 2020).
-All these models consist of an encoder made of multi-layer CNNs followed by one classifier.
+For image classification datasets, we select four DNNs commonly used in mobile/edge-oriented literature: (1) EfficientNet-v2 tan2021efficientnetv2, (2) MobileNet-v3 howard2019searching, (3) MNASNet tan2019mnasnet, and (4) ShuffleNet ma2018shufflenet, with 5.3M, 2.2M, 2.5M, and 1.4M number of parameters respectively.
+For human-activity recognition datasets, we borrow the ConvNet architecture proposed in chang2020systematic.
+All these models consist of an encoder made of multi-layer CNNs followed by a classifier.
 FL Settings.
-We use Flower (Beutel et al., 2020), a customizable open-source FL framework (Python v3.7, Ray v1.11, and Torch v1.12).
+We use Flower beutel2020flower, a customizable open-source FL framework (Python v3.7, Ray v1.11, and Torch v1.12).
 We run all simulations on a server with 80 Intel Xeon(R) E5-2698 CPUs, 8 Tesla V100 GPUs (16GB), and 504GB system RAM.
@@
 We randomly sample 10% of the clients for training in each round in our initial experiments and later show that our results hold when the client participation ratio is gradually increased to 100%.
-Data ($D$) are partitioned using Latent Dirichlet Allocation (LDA) (Blei et al., 2003; Beutel et al., 2022) without resampling (LDA-Alpha=1000).
-We refer to prior research (Blei et al., 2003) about details of LDA’s generative process.
+Data ($D$) are partitioned using Latent Dirichlet Allocation (LDA) blei2003latent; beutel2022flower without resampling (LDA-Alpha=1000).
+We refer to prior research blei2003latent about details of LDA’s generative process.
 After such partition, each client owns $\nicefrac{{|D|}}{{100}}$ local samples.
-Figure 5. Test accuracy when using different values for parameters $\alpha$, $\beta$, $\gamma$ in data selection.
+For each client, we randomly select half of the samples and use them for local training of the model.
+For both APs and UCDs, we consider by default the number of epochs to be 3 with a batch size of 64.
+The other half of the samples are available with more offline time to simulate the further collection of extra data.
+APs and UCDs also train more epochs on the extra data while offline.
+In this setup, we define every unit offline time contributing to one more epoch of training on $20\%$ of these extra local samples.
+To set the hyperparameters for Eqs (2)-(3), we run an experiment on CIFAR-10 with MobileNet-V3.
+We evaluate the trade-off between test accuracy and the workload imposed by data selection on both UCDs and APs.
+This experiment can be executed by the server as a preparatory step before deploying Centaur, using small representative datasets that approximate the clients’ data distribution.
+The results in Figure 4 show that performing data selection reduces latency and energy consumption on UCDs by over 80%.
+Although this approach increases the workload on APs, this is acceptable given their higher computational capacity, since we observe a substantial improvement in model accuracy.
+In our experiments, we set hyperparameter $\alpha,\beta,\gamma$ as $5,3,0$, respectively, as this configuration represents the fair balance between accuracy and resource efficiency.
 […]
@@
 (1) AP Training.
-In this baseline approach we consider UCDs as data collection apparatuses only and therefore, UCDs upload collected and stored data samples to their connected APs, so that the training only happens on APs.
+We consider UCDs as data collection apparatuses only and therefore, UCDs upload collected and stored data samples to their connected APs, so that the training only happens on APs.
 In this case, the complete model (both the encoder and classifier) is updated because APs in general are considered to have sufficient resources.
@@
 In our case, the idea of the mobility model signifies how the mobility patterns of the UCDs (or users) impact the connectivity of the UCD to its AP and eventually to the internet.
-We define an exclusive-Online Association Matrix $\Omega$, which is a binary matrix representing the user’s exclusive location at a given time instance.
+We define an exclusive Online Association Matrix $\Omega$, which is a binary matrix representing the user’s exclusive location at a given time instance.
 Mathematically, $\Omega$ is a binary matrix $\mathcal{T}\times\mathcal{S}$ where $\mathcal{T}$ and $\mathcal{S}$ represent the temporal and spatial granularity of $\Omega$ marix, respectively.
@@
 More specifically, for evaluating Centaur during mobility, we generate the global connectivity matrix $\lambda$ and the generate matrix $\Omega$, both chosen uniformly at random, for each user to simulate the mobility scenario.
-6. Evaluation Results
-We present the results when running Centaur compared to conventional FL training methods.
-We also analyze the impact of data/participation heterogeneity and spatiotemporal coverage on training methods.
 Figure 5.
@@
 Dash lines ( ) above depict the upper-bound accuracy when UCD devices have no resource and connectivity constraints.
+6. Evaluation Results
+We present the results when running Centaur compared to conventional FL training methods and compare the approaches using different metrics.
+We also analyze the impact of data/participation heterogeneity and spatiotemporal coverage on training methods.
+6.1. Metrics
+In addition to the accuracy of the trained global model on a held-out test set we evaluate the performance of our FL framework, compared to other FL alternatives, using the following metrics.
+1) Accuracy.
+This is the classification accuracy of the trained model on a test set hosted by the central server.
+For a model obtained at the end of each training round, the test accuracy is computed as
+$acc\equiv\frac{1}{T}\sum_{i=1}^{T}\mathbb{1}\big(\operatorname*{argmax}\big(f_{E}\circ f_{C}(\mathbf{X}_{i})\big)=\mathbf{y}_{i}\big),$
+where test set has $T$ pairs of $(\mathbf{X}_{i},\mathbf{y}_{i})$ and $\mathbb{1}(\mathrm{C})$ denote the indicator function that outputs 1 if condition $\mathrm{C}$ holds.
+[…]
+Figure 6. Centaur performance vs. other baselines for different sizes of classifiers: small, medium, and large. (Top) image classification, and (bottom) HAR datasets.
+3) Bandwidth.
+As the model size (i.e., the encoder or the classifier) can be different in each FL round, we use fvcore to count the number of parameters that are communicated in each round.
+Based on the model size and number of communications among UCDs, APs, and the server, we compute the amount of bandwidth that is consumed.
+Also, we count the number of sample points that are uploaded from UCDs to APs.
+Figure 7.
+Computation workload of model training on usage-constrained devices (Top Left) and access points (Top Right), and Communication workload of model/data sharing of usage-constrained devices (Bottom Left) and access points (Bottom Right) when achieving specific test accuracy, reported on MobileNetv3 and CIFAR10.
+4) Latency.
+The latency usually has a linear relation with MAC operations, due to the lack of specialized accelerators liberis2021munas; lin2022device.
+Thus, we estimate the latency of model training based on the processor’s frequency as - $\text{Latency}=c*\frac{\text{Total MACs}}{\text{Processor Frequency}}$.
+The ratio between MAC operations and the processor’s instructions, $c$, is typically between $1$ and $2$ based on specific instruction sets/compilers.
+For simplicity, we assume that each MAC operation translates to two instructions in an MCU (i.e., $c=2$).
+Notice that existing processors (e.g., Intel’s Load Effective Address) complete one MAC in one instruction x86assembly), and this only scales the experimental results and does not change the conclusions.
+Finally, communication latency is estimated based on the total amount of data needed to be transmitted, divided by the up-link speed or the down-link speed of devices.
+5) Energy.
+This is the total execution time multiplied by the processor’s consumed power per unit time.
+The energy consumption of communication can also be calculated based on the total time of transmitting data multiplied by the transmitters’ power per unit time.
+We set the values for UCDs and APs in our simulations as follows.
+We assume a typical AP to have a CPU frequency of 2GHz, storage capacity of 4GB, power capacity of 1.5mW/MHz with an uplink speed of 10Mbit/s, downlink speed 100Mbit/s, and communication energy of 10W.
+For a typical UCDs, we assume a CPU frequency of 100MHz, storage capacity of 5MB, power capacity of 0.05mW/MHz with an uplink speed of 2Mbit/s, downlink speed of 2Mbit/s, and communication energy of 0.0001W.
 6.2. Model Accuracy
@@
 We examine three classifiers: small that is only one fully connected (FC) layer of size $z$ (number of classes), medium that has two FC layers of size 64 and $z$, and large that has two FC layers of size 128 and $z$.
-Centaur outperforms AP training and UCD training for different encoders.
 We permute the four encoders and three classifiers and report the test accuracy of Centaur compared to the two other baselines: AP training and UCD training.
@@
 There are several settings in which all FL training methods cannot reach a good accuracy, e.g., NASNet and ShuffleNet on CIFAR100, probably because of their relatively small model sizes compared to data complexity.
-Among them, Centaur still achieves better performance.
-Centaur outperforms UCD training for different classifiers.
+Centaur still achieves better performance compared to the baselines.
 In Figure 6 we present the accuracy of different classifiers for each encoder.
@@
 The test accuracy also tends to be similar across small, medium, and large classifiers.
-In addition, the classifier’s sizes may have less impact on more sophisticated encoders (e.g., EfficientNet and MobileNet), which is also observed in a previous work (Mo et al., 2021).
+In addition, the classifier’s sizes may have less impact on more sophisticated encoders (e.g., EfficientNet and MobileNet), which is also observed in a previous work mo2021ppfl.
 This may be because an appropriate encoder already produces high-quality features for unseen data that are easy to learn (e.g., CIFAR10), and in such a case, classifier sizes do not make any difference in test accuracy.
@@
 This lack of pre-trained encoders can also be the reason that, in HAR datasets, the performance of Centaur and AP are almost the same.
-Figure 7. Centaur performance vs. other baselines for different sizes of classifiers: small, medium, and large. (Top) image classification, and (bottom) HAR datasets.
-Figure 8.
-Computation workload of model training on usage-constrained devices (Top Left) and access points (Top Right), and Communication workload of model/data sharing of usage-constrained devices (Bottom Left) and access points (Bottom Right) when achieving specific test accuracy, reported on MobileNetv3 and CIFAR10.
+These results emphasize that with a properly pre-trained backbone, continual re-training of the encoder on APs is neither necessary nor desirable.
+In transfer learning, it is standard to fix, or only occasionally fine-tune, the encoder and update the lightweight classifier head to adapt to local data.
+In Centaur, we follow this principle: UCDs perform frequent classifier-only updates to rapidly personalize to local data, while APs periodically update the encoder to correct drift in feature extraction.
+This approach achieves competitive accuracy with lower communication and computation costs: frequent classifier updates capture personalization and mobility, while periodic encoder updates maintain convergence and generalization.
+Future work can explore how varying update frequencies affect convergence trade-offs.
+Figure 8. Cost linkage analysis among MAC operations, latency, and energy consumption.
 6.3. Efficiency
 Cost-Accuracy Trade-off.
-We use CIFAR10, and fix MobileNet-v3 as the encoder and medium size for the classifier.
-We compute the test accuracy for a range of MAC, Bandwidth, Latency, and Energy budgets.
+Here, with CIFAR10, MobileNet-v3 as the encoder, and a medium size for the classifier, we compute the test accuracy for a range of MAC, Bandwidth, Latency, and Energy budgets.
 The average size of each sample of CIFAR10 is $30KB$.
-As shown in Figure 8 (top), on both UCD and AP and for all ranges of accuracy, Centaur achieves lower MAC and Latency than both UCD training and AP training.
-In addition, Figure 8 (bottom) shows that communication Bandwidth and Latency of Centaur is almost the same as AP training; however Centaur can achieve a higher test accuracy along with more training steps while consuming more resources.
-Centaur causes much more communication cost than UCD training, because the training only involves the transmission of the classifier, which is much lighter than the encoder and data samples transmitted by Centaur.
+Figure 7 (top) shows on both UCD and AP and across accuracy levels, Centaur achieves lower MAC and Latency than both UCD training and AP training.
+Figure 7 (bottom) shows that communication Bandwidth and Latency of Centaur is almost the same as AP training; however, Centaur can achieve a higher test accuracy along with more training steps when consuming more resources.
+Centaur uses much more communication resources than UCD training, as Centaur transmits both the classifier, although it is much lighter than the encoder, and data samples.
 However, the accuracy of UCD training cannot go further than $83.10\%$, while Centaur can reach up to $89.90\%$ accuracy.
+More reduction in uplink data transmission of UCDs can be achieved by a more advanced data selection strategy, lowering the communication burden on constrained devices.
+This also leads to a notable decrease in overall communication cost compared to AP training, as less data is exchanged between UCDs and APs while improving model accuracy.
 We remark that model training consumes much more energy and causes much more latency than communication.
+Figure 9.
+The impacts of imbalanced data partitioning on test accuracy of AP training, UCD training, and Centaur, on MobileNetv3 with CIFAR10 (annotated texts are examples of the classes present on a UCD)
+Figure 10. The impacts of imbalanced data partitioning on test accuracy of AP training, UCD training, and Centaur, on ConvNet and UCIHAR dataset dataset.
 Correlation among Cost Metrics.
@@
 Such linear relationships are because of the assumption we made when computing Latency and Energy; however, the actual cost and the relationship among these metrics might not deviate much in practice.
-Based on the above-mentioned correlation, we can add Latency (or Energy) of model training as well as the communication Bandwidth to get the overall workload.
+Based on such correlations, we can add Latency (or Energy) of model training as well as the communication Bandwidth to get the overall workload.
 The results shown in Figure 1 indicate that the workload still follows a similar pattern as the computation in model training, because the training workload significantly overweight the communication workload.
@@
 Centaur is robust to data heterogeneity.
-To create imbalanced non-IID data partitions among clients, we use LDA as defined in § 4.
+To create imbalanced non-IID data partitions among clients, we use LDA as defined in § 6.1.
 Next, we set different values for LDA-Alpha to manipulate the levels of non-IID data partitions.
 Figure 10 and Figure 10 show the results when LDA-Alpha are changing from $0.001$, to $0.1$, to $1$, and to $1000$.
-The smaller is LDA-Alpha, the less balanced is the dataset.
-We also annotate some examples of class distribution in Figure 11 demonstrating that LDA-Alpha$=0.001$ generates almost $1$ class per client, while LDA-Alpha$=1000$ generates almost uniform classes per client.
+The smaller LDA-Alpha is, the less balanced the dataset is.
+We also annotate some examples of class distribution in Figure 10, demonstrating that LDA-Alpha=0.001=0.001 generates almost $1$ class per client, while LDA-Alpha=1000=1000 generates almost uniform classes per client.
 Results show that in all cases Centaur can reach higher test accuracy than both AP and UCD training.
-We remark that due to the retraining of the encoder, both Centaur and AP training can cause more fluctuations than UCD training; this is especially obvious when data distribution tends to be IID (e.g., when LDA-Alpha$=1000$ in Figure 11).
+We remark that due to the retraining of the encoder, both Centaur and AP training can cause more fluctuations than UCD training; this is obvious when data distribution tends to be IID (e.g., when LDA-Alpha=1000=1000 in Figure 10).
 Centaur scales well with higher total and participating clients.
-We further scale the FL training problem with the number of total clients from 10 to 1000.
+Here, we scale the FL training problem with the number of total clients from 10 to 1000.
 $10\%$, $20\%$, and $50\%$ of them are selected respectively as participating clients in each round.
@@
 Also, with such more fragmented data partitioning, Centaur has much accuracy gain when compared to UCD training.
-Figure 9. Cost linkage analysis among MAC operations, latency, and energy consumption.
-Figure 10.
-The impacts of imbalanced data partitioning on the test accuracy of AP training, UCD training, and Centaur, reported on MobileNetv3 on CIFAR10 dataset (annotated texts give examples of the classes present on the UCDs)
-Figure 10.
-The impacts of imbalanced data partitioning on the test accuracy of AP training, UCD training, and Centaur, reported on MobileNetv3 on CIFAR10 dataset (annotated texts give examples of the classes present on the UCDs)
-Figure 11. Impacts of imbalanced data partitioning on the test accuracy of AP training, UCD training, and Centaur, reported on ConvNet and UCIHAR dataset dataset.
+This experiment focuses on test accuracy and does not account for the orchestration costs of running Centaur in dynamic mobile environments.
+However, such scalability challenges are not unique to Centaur but are inherent to any FL framework.
+We also note that practical deployment at larger scales necessitates additional server-side coordination.
+Centaur is designed to be modular, enabling integration into existing FL stacks.
+Nevertheless, evaluating the end-to-end overhead of managing varying participation rates and asynchronous aggregation remains open.
 6.5. Performance of Centaur under Mobility
-We evaluate situations when the connection probability $\lambda$ changes based on the mobility model defined in §4.
-For a more readable visualization, we bucket $\lambda$ in three different ranges of $\lambda\in$ [0.1,0.4], $\lambda\in$ [0.4,0.7], and $\lambda\in$ [0.7,1.0] and report the corresponding Accuracy of Centaur when the UCDs are mobile.
-Results reported in Figure 13 show that Centaur always has higher efficiency than standard UCD training while incurring a higher latency than direct AP training.
-Specifically, regarding the much important UCD side, Centaur gains $6.45\%$, $5.64\%$, and $4.11\%$ higher accuracy with $51.73\%$, $60.48\%$, and $68.35\%$ lower cost both in terms of energy and latency across the different ranges of $\lambda$.
-Interestingly, we also observe a rise in accuracy with lower connectivity probability.
-The primary reason is attributed to the overall design of our experimental setup whereby we make a realistic assumption that when mobile with limited connectivity, these UCDs can gather more data and perform the data selection.
-Additionally, with this judiciously selected data, the device trains the classifier for an extended number of epochs thus boosting its accuracy.
-7. Discussion and Related Work
-Multitier FL.
-Compared to the default client-server FL, client-edge-server FL (Liu et al., 2020; Abad et al., 2020) considers a middle layer, e.g., a cellular base station, to orchestrate FL training across clients and a central server.
-In such a setup, clients in the last tier participate with several devices (Cho et al., 2022; Diao et al., 2020; Zhang et al., 2021).
-This is shown to achieve better trade-offs between accuracy and bandwidth consumption.
-However, these works assume local training of the whole model on the client devices and do not utilize the capabilities of a resourceful edge device for aiding a constrained device during training, as they are only used for aggregation/communication.
-This makes it impractical in scenarios involving UCDs.
-On-device Training.
-Most prior works only consider the challenges of on-device inference (i.e., forward pass) (Fedorov et al., 2019; Banbury et al., 2021; Xu et al., 2022; Liberis et al., 2021), or work on sparse training (Li et al., 2021a; Bibikar et al., 2022).
-Sparse training or pruning still requires computational resources (especially memory) that UCDs do not have.
-Few works propose solutions for full on-device training (i.e., forward and backward pass) (Lin et al., 2022), but not for FL scenarios as they often make impractical assumptions like the availability of large datasets for network architecture search (Liberis et al., 2021) or unconstrained memory for hosting the full model (Diao et al., 2020; Zhang et al., 2021).
-Transfer learning can be used on UCDs to fix the encoder architecture, which in turn restricts the extraction of newer features, and only trains the last layer for the personalized training (Zhang et al., 2021).
-Regarding complex tasks, training the fully functional model on UCDs is challenging, because the on-device model still needs the capability of learning all features by itself (Li et al., 2021a; Bibikar et al., 2022), and the computational resources at the upstream devices are unutilized.
-Figure 12. The impact of participating clients and total clients on test accuracy reported on MobileNetv3 and CIFAR10.
-Communication/Power Efficiency.
-Using a pre-trained encoder without further training might fail to capture discerning features from the newly generated data, causing lower performance for the ultimate classifiers.
-On the other hand, passing all collected sensor data from UCDs to APs (even with data compression) is not always feasible and it might consume a lot of energy and bandwidth.
-Importantly, network availability for UCDs can be fragile due to the low-energy wireless communication protocols they use and the mobility of these UCDs.
-As UCDs are becoming stand-alone, it is expected that they can operate as independent devices without internet or permanent connection to an upstream device, Centaur takes a major step in unlocking more functionalities of such development.
-Compression techniques (Lim et al., 2019) or topology-aware hierarchical aggregation (Luo et al., 2020) can reduce Latency by taking advantage of the locality in networks and continuously balancing workloads across clients.
-However, compression and hierarchical aggregation alone do not allow for seamlessly incorporating UCDs in FL.
-Our proposed method comprises different methods to offer customizable trade-offs between test accuracy and communication cost in data- and topology-agnostic manners, while allowing additional use of the recent compression and hierarchical aggregation techniques for further efficiency.
-Training on UCDs.
-By default, FL algorithms are not specifically designed for a multi-device environment (Lee et al., 2019; Islam and Nirjon, 2019).
-Training DNNs on tiny MCU devices is extremely challenging due to limited memory, computation, and power resources.
-There are recent efforts to enable running ML models on a micro-controller with highly limited memories, such as TensorFlow Lite Micro (David et al., 2021) or Tiny Training Engine (Lin et al., 2022).
-Such developments, usually based on a combination of hardware and software novelties, can be easily integrated into Centaur to provide further efficiency.
-As we observe, a better training framework can not only enable deploying larger classifiers on UCDs, but can also allow us to calculate more gradients for better data selection.
-While it is difficult to train large DNNs on UCDs, these devices are typically parts of a personal network, and therefore, they can privately communicate/coordinate with other surrounding devices.
-Our work considers such a hierarchy for unlocking more utilities while providing better efficiency.
+Figure 11. The impact of participating vs. total clients on test accuracy for MobileNetv3 with CIFAR10.
 Figure 12.
 The performance of access point (AP) training, usage-constrained device (UCD) training, and partition-based federated learning (Centaur) under different connection probabilities $\lambda$, reported on MobileNetv3 and CIFAR10.
-8. Limitations and Future Work
-First, we accommodate FedAvg (McMahan et al., 2017b) in our current implementation, but Centaur is compatible with other FL algorithms, e.g., FedProx (Li et al., 2020), FedAdam (Reddi et al., 2020), or DM-PFL (Zhang et al., 2023).
-On the other hand, layer-wise FL methods, such as FedMA (Wang et al., 2020), may have interesting interactions as they also conduct model training in a partitioned fashion.
-Second, more algorithmic changes can be done to existing AP training, but we emphasize that whatever we do for AP training can only be useful for applications where all data is available at AP.
-In many emerging applications, UCDs are the main devices that collect and process data in the first place, and we do not want UCDs to send all the data to AP and put the burden of data selection on the APs shoulders.
-Third, we do not perform model selection or architecture search, and our model partitioning simply follows the encoder-classifier style of benchmark networks, but one can enhance the model initialization and partitioning with state-of-the-art advancements in neural architecture search (Li et al., 2021b; Huang et al., 2023).
-Fourth, our goal is to present a proof of concept for a multi-device FL that encompasses UCDs and APs.
-We present a real testbed and perform actual measurements, but in the rest of our experiments, we utilize GPU servers for running end-to-end FL scenarios.
-The constraints of UCDs are simulated according to the size and type of the model they can train.
-As a topic for future investigation, and to achieve more precise calculations of compute and storage resources, the end-to-end process of Centaur could be implemented on real-world UCDs and APs.
-Fifth, we considered a one-to-one relationship between UCDs and APs, but one can expand Centaur into a many-to-one setting.
-This, however, needs tackling the challenge of aggregating diverse classifiers trained by various sensing devices.
-Sixth, we evaluated Centaur on supervised FL tasks, but it is possible to suggest a supplementary approach for implementing semi-supervised or unsupervised tasks.
-Loss-based data selection has been studied very well in the ML community and shows promising advantages.
-Finally, dealing with outliers and distribution shifts are important challenges but we believe these are topics outside the scope of this paper.
+We study when the connection probability $\lambda$ changes based on the mobility model defined in §6.1.
+For a more readable visualization, we bucket $\lambda$ in three different ranges of $\lambda\in$ [0.1,0.4], $\lambda\in$ [0.4,0.7], and $\lambda\in$ [0.7,1.0] and report the corresponding accuracy when the UCDs are mobile.
+Figure 12 shows Centaur always has higher efficiency than standard UCD training while incurring a higher latency than direct AP training.
+Specifically, Centaur gains $\%, 5.645.64\%$, and $4.11\%$ higher accuracy with $51.73\%$, $60.48\%$, and $68.35\%$ lower cost both in terms of energy and latency across the different ranges of $\lambda$.
+Interestingly, we also observe a rise in accuracy with lower connectivity probability.
+The main reason is the design of our experimental setup: we realistically assume that with limited connectivity, UCDs devices can collect more data and perform data selection locally.
+Using this selected data, each device then trains the classifier for more epochs, which further improves its accuracy.
+7. Discussion
+While Centaur shows strong empirical performance and feasibility on constrained devices, some of our assumptions need further study.
+We outline key limitations, and opportunities for future work, below.
+(1) Centaur assumes each UCD has a trusted, resourceful personal AP (e.g., a smartphone or router) that is always available.
+This simplifies privacy and connectivity but may not hold in transient or multi-user settings.
+(2) We used RaspberryPi as a reproducible proxy for constrained devices, though it is more capable than many real UCDs.
+Results may thus overestimate feasibility; testing on truly constrained hardware is needed to validate our claims.
+(3) Freezing the encoder and training only the classifier reduces on-device compute and memory but limits adaptability: if local data diverge from the pretrained one, classifier-only updates may be insufficient.
+Studying trade-offs in encoder vs.
+classifier update frequency remains open.
+(4) Our on-device selection uses loss and gradient norm for importance sampling, but this remains heuristic without formal convergence or variance guarantees.
+It may also be sensitive to outliers, mislabeled, or out-of-distribution data, potentially biasing updates.
+(5) Centaur reduces on-device workload but increases AP computation and network use.
+We did not quantify orchestration and scaling costs with many asynchronous users.
+Although fewer samples are sent than baselines, uplink and energy costs for constrained devices and overall communication efficiency need further analysis.
+(6) We use FedAvg in our implementation, but Centaur can accommodate other FL aggregation strategies.
+Layer-wise FL methods that perform model training in a partitioned manner, such as FedMA wang2020federated, can also be integrated.
+Also, we evaluated Centaur on supervised tasks, yet a complementary approach can be suggested for unsupervised tasks.
 8. Conclusions
-To include constrained devices in an FL task, we offer an efficient training approach that utilizes data selection to enhance learning while minimizing computational costs.
-Additionally, we employ partition-based training and aggregation, wherein constrained devices train a specific part of the model, aided by more capable devices that train the complete model.
-Our assessments validate the benefits of our proposed FL framework, Centaur, in enhancing model accuracy while simultaneously reducing system overhead.
-Furthermore, we propose several promising directions for future studies in the area of efficient FL on constrained sensory devices.
+We propose an efficient training method that leverages data selection to improve learning and reduce computational costs in multidevice federated learning.
+Our approach allows partition-based training and aggregation on resource-constrained devices, supported by resourceful companion devices.
+Our evaluations over various scenarios confirm that our proposed solution improves model accuracy while reducing system overhead.